### v1

# Unveiling the Mathematical Reasoning in DeepSeek Models: A Comparative Study of Large Language Models

Abstract
With the rapid evolution of Artificial Intelligence (AI), Large Language Models (LLMs) have reshaped the frontiers of various fields, spanning healthcare, public health, engineering, science, agriculture, education, arts, humanities, and mathematical reasoning. Among these advancements, DeepSeek models have emerged as noteworthy contenders, demonstrating promising capabilities that set them apart from their peers. While previous studies have conducted comparative analyses of LLMs, few have delivered a comprehensive evaluation of mathematical reasoning across a broad spectrum of LLMs. In this work, we aim to bridge this gap by conducting an in-depth comparative study, focusing on the strengths and limitations of DeepSeek models in relation to their leading counterparts. In particular, our study systematically evaluates the mathematical reasoning performance of two DeepSeek models alongside five prominent LLMs across three independent benchmark datasets. The findings reveal several key insights: 1). DeepSeek-R1 consistently achieved the highest accuracy on two of the three datasets, demonstrating strong mathematical reasoning capabilities. 2). The distilled variant of LLMs significantly underperformed compared to its peers, highlighting potential drawbacks in using distillation techniques. 3). In terms of response time, Gemini 2.0 Flash demonstrated the fastest processing speed, outperforming other models in efficiency, which is a crucial factor for real-time applications. Beyond these quantitative assessments, we delve into how architecture, training, and optimization impact LLMs’ mathematical reasoning. Moreover, our study goes beyond mere performance comparison by identifying key areas for future advancements in LLM-driven mathematical reasoning. This research enhances our understanding of LLMs’ mathematical reasoning and lays the groundwork for future advancements.
1 Introduction
As Artificial Intelligence (AI) continues to advance at an unprecedented pace, an array of powerful Large Language Models (LLMs) has emerged, including OpenAI’s GPT-4o, o1, o3, Claude 3.5, Llama 3.2, Qwen 2.5, and Gemini 2.0 [1, 12, 28, 30, 48, 57, 69, 82, 102]. Driven by cutting-edge advancements in Deep Neural Networks (DNNs), these models integrate human cognition elements to enhance problem-solving and decision-making [49]. Serving as a guiding light in natural language processing (NLP) [65], healthcare [21, 60, 61], clinical textual processing [22, 59, 104], biomedical image analyses [60], code generation [58], decision support [51], multimodal data analytics [83, 90], and mathematical reasoning [2, 52, 72, 76, 91, 103, 107], LLMs push the boundaries of AI capabilities, approximating human-like reasoning through sophisticated statistical inference [1, 48, 69]. However, despite their transformative potential, these models face notable limitations. Their high computational demands pose significant barriers to broader accessibility, making large-scale implementation costly [69]. Moreover, while LLMs perform well in general contexts, they often struggle with specialized tasks, exhibiting inconsistencies in performance. Multimodal models, for example, continue to face challenges in spatial reasoning and real-world physics, while AI-assisted code generation frequently produces syntactically correct yet functionally flawed outputs, requiring human oversight [58, 103]. These constraints underscore the ongoing need for refinement and innovation in AI research to bridge the gap between artificial and human intelligence.
In particular, GPT-4o, released by OpenAI in May 2024, is a multimodal model capable of processing text, images, and voice with remarkable efficiency. Leveraging an advanced transformer architecture, it surpasses GPT-3 in critical areas such as mathematical reasoning and language comprehension [1, 23, 44, 75]. With an estimated 2 trillion parameters, GPT-4o is significantly larger than its predecessors, enabling substantial improvements in performance and adaptability. Meanwhile, other cutting-edge models, such as o1 and o3, have been introduced in 2024 and 2025, respectively. In particular, o1 enhances reasoning capabilities by incorporating a Chain-of-Thought (CoT) approach [80], where it renders intermediate reasoning steps before reaching a final answer closely mirroring the way humans process complex problems [103]. More significantly, o3 advances reasoning performance through a "simulated reasoning" process, in which the model actively generates and evaluates multiple solution paths [6]. By pausing, reflecting, and adjusting its approach before delivering a final answer, o3 exhibits human-like reasoning, allowing for more nuanced and adaptive problem-solving, particularly in highly complex scenarios [67]. However, none of the GPT models, including GPT-4o, o1, and o3, are open-source, as they remain proprietary.
In addition, Claude 3.5, released in 2024, is built on previous versions. It emphasizes safety, alignment, and performance, with improvements in reasoning, language understanding, and handling complex tasks like text and code generation [30]. With 250B parameters, it surpasses earlier models in accuracy and ethical alignment. It supports up to 200K tokens for extended context, enabling better processing of larger inputs. Notably, enhanced by reinforcement learning from human feedback (RLHF) [41, 85] and Constitutional AI, it reduces undesirable responses, biases, and better aligns with human intent. Claude 3.5 excels in specialized areas like coding and scientific reasoning, with improved transparency and ethical safeguards [30, 69].
Furthermore, Llama-3.3, presented in 2024, is the latest version of LLM in Meta AI family, following Llama-1 and Llama-2. Llama-3.3 advances further with 70B parameters and 128K token context window, improved by grouped-query attention for better efficiency [28, 63]. Llama 3.1 excels in coding, logical problem solving, and low-resource language tasks. Unlike closed models such as GPT series, it remains open-weight and freely accessible for research and commercial use, but is restricted to text-only input [26, 63, 69]. Safety measures, such as automated red-teaming and filtered training data, help minimize undesirable outputs. Moreover, Gemini 2.0 Flash is the latest multimodal LLM of Google, building on versions 1.0 and 1.5 to offer more robust generative AI capabilities across text, images, audio, and video.
Moreover, Gemini 2.0 Flash, initially introduced as an experimental variant, provides significant speed and efficiency gains over its predecessor, Gemini 1.5 Flash, without sacrificing the efficiency [43, 45, 82]. Notably, it outperforms Gemini 1.5 Pro on key benchmarks while operating at twice the speed. It enables the incorporation of agentic AI and native use, allowing the model to call external functions (Google Search and Maps) and integrate streaming data for expanded real-time applications. By combining better performance in tasks such as math, code generation, and multilingual audio output with enhanced efficiency. Gemini 2.0 aims to deliver comprehensive, cost-effective AI solutions for both developers and end users [43, 82]. Lastly, Qwen2.5, released in September 2024, is the latest iteration in the Qwen series, following Qwen2 in June 2024 and the original Qwen in August 2023. Qwen1.5 featured models up to 72B parameters, emphasizing efficiency and open-source accessibility. Qwen2 introduced improved reasoning, multilingual support, and coding capabilities, with models scaling up to 72.71B parameters. [94, 106]
Besides, established in 2023 as a research initiative to push the boundaries of artificial general intelligence (AGI), DeepSeek models set out to overcome existing limitations by developing specialized models focused on efficiency, adaptability, and domain expertise [8, 20, 38, 37, 57, 48, 56, 76]. In 2024, the Mixture-of-Experts (MoE), an efficiency-driven architecture that leverages sparse activation to reduce computational overhead [38, 56, 57], for DeepSeek was introduced. This was followed by the launch of DeepSeek Coder, a suite of code-focused models ranging from 1B to 33B parameters, designed to streamline software development workflows. Meanwhile, DeepSeek Math, trained on 120B math-related tokens, was developed to handle advanced mathematical and symbolic reasoning tasks [76]. Expanding its model portfolio, DeepSeek introduced the V2 and V3 series [56, 57]. V2 implemented multi-head latent attention (MLA) alongside a MoE system with 236B parameters in total, of which only 21B were active per query, optimizing computational efficiency [56]. V3, an open-source model, further enhanced efficiency with 671B total parameters, activating only 37B per query, excelling in complex reasoning tasks while minimizing resource demands and reliance on supervised data [57]. In 2025, DeepSeek made a significant breakthrough with R1 Zero, incorporating self-verification, reflection, and extended CoTs. Notably, DeepSeek-R1 was recently presented, specifically designed for mathematical, coding, and logical problem-solving, enhancing autonomous decision-making and precision in both research and enterprise applications [37, 38, 76]. To extend accessibility, an open-sourced suite of distilled models for DeepSeek, optimized for deployment in resource-constrained environments such as edge computing platforms and low-memory systems. These models preserve scalability and cost-effectiveness, making cutting-edge AI more accessible across diverse applications.
Notably, mathematical reasoning usually poses intricate challenges without straightforward solutions [3, 55]. Unlike routine tasks that follow established frameworks, mathematical reasoning demands creativity, abstract thinking, and advanced cognitive skills. In this study, we systematically evaluate the mathematical reasoning performance of a wide array of LLMs, with a particular emphasis on DeepSeek models, to further unveil machine creativity in mathematical reasoning. Leveraging multiple independent datasets and quantitative metrics, we conduct a comprehensive empirical analysis to assess and compare their capabilities. Our findings provide a detailed evaluation of prominent LLMs in mathematical reasoning while highlighting the advancements and unique strengths of DeepSeek models.
2 Related Works
Recent research has increasingly focused on summarizing and evaluating the performance of LLMs in problem-solving and code generation, areas where general-purpose LLMs excel in text-based tasks but often struggle with mathematical precision and structured reasoning [24, 36, 48, 77, 86, 97]. To address these limitations, LLMs have prioritized enhancing reasoning capabilities and improving computational efficiency in next-generation models, aiming to bridge the gap between linguistic fluency and robust problem-solving skills.
Specifically, in 2024, Wang et al., summarized and detailed the specifications of Multimodal Large Language Models (MLLMs), standing out at the forefront of AI [83]. In brief, Wang et al., overviewed a wide array of data modalities, such as text, images, audio, and sequential data [83]. In fact, MLLMs play a crucial role in multimodal understanding tasks by integrating text and image information to achieve more intelligent and comprehensive understanding and reasoning. In this field, Wang et al., outlined the developments, advancements, and utilization of multiple MLLMs, including MiniGPT-4, InstructBLIP, and Wiki-LLaVA are three highly regarded models, along with other related MLLMs such as 3DMIT, GroundingGPT, ModaVerse, Vary-toy, LLaVAMOLE, and CogCom [83].
In addition, at the end of 2024, Zhong et al. established a comprehensive and extensive evaluation of o1-preview (the early version of o1) to showcase the advancement of OpenAI’s o1-preview across a diverse array of complex reasoning tasks, spanning multiple domains, including computer science, mathematics, natural sciences, medicine, linguistics, and social sciences [103]. Through extensive testing, o1-preview highlighted remarkable capabilities, often achieving human-level or superior performance in areas ranging from coding challenges to scientific reasoning and from language processing to creative problem-solving [103]. Key findings include: 1). 83.3% success rate in solving complex competitive programming problems, surpassing human experts. 2). Superior ability in generating coherent and accurate radiology reports, outperforming other evaluated models. 3). 100% accuracy in high school-level mathematical reasoning tasks, providing detailed step-by-step solutions. 4). Advanced natural language inference capabilities across general and specialized domains like medicine. 5). Impressive performance in chip design tasks, outperforming specialized models in areas such as script generation and bug analysis. 6). Remarkable proficiency in anthropology and geology, demonstrating deep understanding and reasoning in these specialized fields. 7). Comprehensive financial knowledge and strong capabilities in quantitative investing. 8). Effective performance in social media analysis, including sentiment analysis and emotion recognition. 9). Excellent performance in educational measurement and psychometrics, demonstrating a solid grasp of standard psychometric concepts that are equivalent to or beyond a first-year master’s or doctoral student’s level.
Furthermore, Yang et al. investigated the capabilities of advanced LLMs, particularly the o1 model, in literary analysis [96]. Given the prestige of the Nobel Prize and its emphasis on cultural, historical, and linguistic depth, applying LLMs to these works offers valuable insights into both human and AI approaches to literary interpretation [96]. The study employed qualitative and quantitative evaluations to assess coherence, creativity, and fidelity to the text, shedding light on the strengths and limitations of AI in domains traditionally dominated by human expertise. While LLMs demonstrated remarkable analytical capabilities, particularly in structured tasks, they struggled with emotional nuance and coherence—areas where human interpretation remains unparalleled. This research underscores the transformative potential of human-AI collaboration in the humanities, paving the way for new opportunities in literary studies, textual analysis, and interdisciplinary research. By leveraging AI’s analytical power alongside human intuition, this study highlights promising directions for AI-assisted literary interpretation and beyond.
Moreover, Xu et al. demonstrated that LLMs have made significant advancements in clinical decision-making, particularly those leveraging in-context demonstrations and specialized medical fine-tuning [92]. These models exhibit strong performance in medical language processing but continue to face challenges in real-time adaptability, multi-step reasoning, and handling complex medical tasks. Agent-based AI systems aim to overcome these limitations by incorporating reasoning traces, contextual tool selection, knowledge retrieval, and both short- and long-term memory. These features enable medical AI agents to manage complex clinical scenarios where decision-making requires real-time interaction with the environment [92]. Unlike conventional model-based approaches that treat medical queries as isolated questions, medical AI agents approach them as dynamic, multi-faceted tasks, allowing them to function more like human doctors. Xu also investigated the selection of the backbone LLM for medical AI agents, which serves as the foundation for their reasoning and action generation. Specifically, Xu et al. [92] examined the capabilities of the emerging o1 model and its impact on agents’ reasoning, tool-use adaptability, and real-time information retrieval across diverse clinical scenarios, including high-stakes environments such as intensive care units (ICUs). The findings highlighted o1’s potential to enhance diagnostic accuracy and consistency, paving the way for more intelligent, responsive AI tools that support improved patient outcomes and more effective decision-making in clinical practice [92].
Notably, Ahn et al. [3] addressed four critical aspects of LLMs in mathematical reasoning: 1). mathematical problem types and datasets, 2). techniques for enhancing LLM performance (such as prompt engineering and fine-tuning), 3). factors affecting model effectiveness (including scale and pre-training data), and 4). challenges such as brittleness and inconsistency, which often cause models to generate different answers for similar problems.
Besides, Liang et al. presented a comprehensive overview on the applications of AI in mathematical research, highlighting the transformative role AI has begun to play in this domain [55]. Traditionally, AI advancements have heavily relied on theoretical foundations provided by mathematics and statistics. However, recent developments in AI, particularly in reinforcement learning (RL) and LLMs, have demonstrated the potential for AI to contribute back to mathematics by offering flexible algorithmic frameworks and powerful inductive reasoning capabilities that support various aspects of mathematical research. This survey aimed to establish a bridge between AI and mathematics, providing insights into the mutual benefits and fostering deeper interdisciplinary understanding. In particular, Liang et al. discussed that while current AI and LLMs may struggle with complex deductive reasoning, their inherent creativity, the capability to generate outputs at high throughput based on recognition of shallow patterns, holds significant potential to support and inspire mathematical research [55]. Furthermore, Liang et al. [55] addressed the lack of cross-disciplinary communication: mathematicians may not fully comprehend the latest advances in AI, while AI researchers frequently prioritize benchmark performance over real-world applications in frontier mathematical research.
Lastly, this work seeks to close that gap, offering a detailed exploration of AI fundamentals, its strengths, and its emerging applications in the mathematical sciences. While various studies have begun exploring in-depth comparisons between DeepSeek-R1 and other leading models [37, 38, 76], few provide a comprehensive analysis of their mathematical reasoning capabilities. Thus, in this work, we introduce a holistic empirical study to systematically compare the mathematical reasoning performance of DeepSeek-R1 with its peer models. Our goal is to contribute to a deeper understanding of DeepSeek-R1’s strengths and limitations while inspiring further advancements in the future works, using a wide array of peer models, datasets, and quantitative metrics.
3 Methodology
In this section, we provide a comprehensive overview of our experimental framework, detailing the datasets utilized, the quantitative evaluation metrics, as well as the overall methodology. We also present a justification for selecting these datasets as benchmarks, emphasizing their relevance, diversity, and suitability in effectively assessing LLM performance in mathematical reasoning. By establishing a consistent evaluation framework, we aim to ensure a thorough and insightful comparison of model capabilities across various reasoning tasks.
Figure 1: This figure illustrates the experimental framework of this work, including five vital components.
3.1 Experimental Frameworks
The framework of this empirical study is illustrated in Figure 1 and consists of five key components: 1). Benchmark Data: We utilized three representative benchmark datasets [16, 39, 62] to evaluate the mathematical reasoning capabilities of multiple LLMs. 2). Zero-Shot Prompting: Zero-shot prompting was employed to design all prompts [42], ensuring a standardized evaluation approach. 3). Language Model Selection: A diverse set of representative LLMs was selected, including Google’s Gemini 2.0 Flash [45], OpenAI’s GPT-4o [1], o1-mini [46], o1 [103], o3-mini [6], DeepSeek’s DeepSeek-R1-Distill-Qwen-1.5B [37], hereafter referred to as DeepSeek-1.5B, and DeepSeek-R1 [56]. 4). Evaluation: Accuracy was used as a critical quantitative metric to assess and compare the mathematical reasoning performance of these models [17, 18, 19]. 5). Comparison Between Models: A comprehensive statistical analysis was conducted to evaluate and contrast the performance of each LLM, offering deeper insights into their mathematical reasoning capabilities. Overall, this structured framework enabled a holistic and systematic comparison of LLMs in mathematical reasoning.
3.2 Datasets
To evaluate the mathematical reasoning capabilities of LLMs, we employ well-established benchmark datasets that encompass diverse mathematical domains, question types, and difficulty levels. The selected datasets include Math Competition (MATH) [40], Grade School Math 8K (GSM8K)  [16], and Massive Multitask Language Understanding (MMLU) [39], covering topics ranging from elementary arithmetic to advanced algebra, logic, and mathematical competitions. From each benchmark, we carefully select a representative subset of problems to ensure a balanced evaluation while maintaining the diversity and complexity of mathematical reasoning tasks. These subsets are chosen for their broad coverage of mathematical domains, variety in question formats, and representation of different complexity levels, ensuring a comprehensive assessment of LLMs’ symbolic manipulation, logical reasoning, and problem-solving abilities [35, 74, 101]. In total, our evaluation spans 2,178 mathematical problems across multiple benchmark datasets, ensuring a comprehensive assessment of LLMs’ mathematical reasoning capabilities.
By evaluating LLMs on these benchmarks, we identified their strengths and limitations in mathematical reasoning, aiding in the development of more robust and capable models. Table 1 provides an overview of the datasets utilized in this study.
Table 1: Overview of Benchmark Data subsets Used for Evaluation
Dataset Category Question Type Domain Number of Problems
MATH Mathematics competitions Free-form numberic Algebra, Geometry, Number Theory, Combinatorics 262
GSM8K Grade School Math Word Problems Arithmetic 1320
MMLU College Mathematics Multiple-choice Algebra, Calculus, Discrete Math 100
Abstract Algebra Multiple-choice Group Theory, Rings, Algebraic Structures 100
Formal Logic Multiple-choice Proofs, Logical Reasoning 126
High School Mathematics Multiple-choice General Mathematics 270
3.2.1 MATH
MATH is a comprehensive benchmark dataset designed to evaluate mathematical reasoning across a wide range of problem-solving scenarios [40]. Unlike conventional mathematics problem datasets, which primarily assess algorithmic proficiency, MATH focuses on advanced reasoning and heuristic-driven problem-solving.
Specifically, MATH is derived from high-level mathematics competitions, including the AMC 10, AMC 12, and AIME [40]. These problems are inherently more complex than standard K-12 mathematics tasks, requiring the application of non-trivial problem-solving strategies, logical inference, and domain-specific heuristics rather than direct formulaic computations. Meanwhile, MATH is widely regarded as a difficult benchmark, with LLMs achieving accuracy rates between 3.0% and 6.9% [40]. Despite these low overall scores, LLMs demonstrate some degree of mathematical competency, as evidenced by: Up to 15% accuracy on the easiest difficulty level [40]. The ability to generate step-by-step solutions that, while sometimes incorrect, remain coherent and contextually relevant. To gauge the dataset’s difficulty, human performance was also considered. For instance, a computer science Ph.D. student with no particular affinity for mathematics achieved approximately 40% accuracy. In addition, a three-time IMO (International Mathematical Olympiad) gold medalist attained 90% accuracy. Thus, these findings indicate that MATH is challenging for both human solvers and LLMs, making it an invaluable dataset for assessing and advancing mathematical reasoning capabilities in both LLMs and human problem solvers.
3.2.2 GSM8K
GSM8K, introduced in 2021, is a benchmark dataset consisting of 8,500 high-quality grade school-level math problems  [16]. Designed to incorporate high linguistic diversity while relying on fundamental mathematical concepts, it presents a unique challenge for state-of-the-art LLMs. Although the underlying math is relatively simple, the diverse problem formulations create significant hurdles, preventing many models from achieving consistently high accuracy. GSM8K is structured into 7,500 training problems and 1,320 testing problems, all carefully crafted by expert human problem writers. The problems primarily involve elementary arithmetic operations and typically require 2 to 8 logical steps to reach a solution. This dataset is widely used to evaluate logical reasoning and mathematical proficiency in LLMs and serves as a benchmark for various assessments, including the LLM Leaderboard. Notably, while some GSM8K problems are conceptually straightforward, they can still be challenging for even the most advanced LLMs, often exhibiting high variability in responses when the same problem is presented in slightly different ways. This highlights the ongoing difficulty of achieving robust mathematical reasoning in language models. In this study, we selected the 1.32K mathematical test set from GSM8K to systematically evaluate and compare the mathematical reasoning capabilities of each LLM.
3.2.3 MMLU
The MMLU dataset is a comprehensive benchmark comprising multiple-choice questions from a diverse range of academic disciplines. Spanning 57 distinct tasks, it covers subjects in the humanities, social sciences, hard sciences, and mathematical reasoning, reflecting the breadth of knowledge that is essential for various fields of study [39]. MMLU’s questions were manually curated by graduate and undergraduate students from publicly available sources, including practice questions from standardized exams such as the Graduate Record Examination (GRE) and the United States Medical Licensing Examination (USMLE). Additionally, it features questions designed for undergraduate courses and readers of Oxford University Press books. Notably, its mathematical reasoning component encompasses multiple subfields, categorized into areas such as "Abstract Algebra," "College Mathematics," "Formal Logic," and "High School Mathematics." For instance, questions in the "Abstract Algebra" category originate from professional mathematical practice, while the "High School Mathematics" category includes problems akin to those found in standard high school exams. In total, MMLU comprises 15,908 questions, divided into three subsets: a). A few-shot development set containing five questions per subject. b). A validation set with 1,540 questions, used for selecting hyperparameters. c). A test set with 14,079 questions, ensuring a rigorous evaluation Each subject within MMLU includes a minimum of 100 test questions, making it a more extensive and challenging assessment than most standard exams. This broad and structured dataset serves as a critical benchmark for evaluating the reasoning and problem-solving capabilities of LLMs.
3.3 Evaluate LLMs in Mathematical Reasoning
As discussed previously, multiple LLMs have demonstrated remarkable abilities in mathematical reasoning [105]. To systematically evaluate the mathematical reasoning capabilities of various LLMs, we conducted experiments using API-based access to a diverse set of models. The selected models include Google’s Gemini-2.0-flash, OpenAI’s GPT-4o, o1-mini, o1, and o3-mini, and DeepSeek’s DeepSeek-1.5B and DeepSeek-R1. These models were accessed through their respective API services using secure API keys from Google, OpenAI, DeepSeek, and Openrouter.
3.3.1 Prompting Strategy
The effectiveness of LLMs in mathematical reasoning is highly dependent on the design of input prompts [5, 10, 25]. To ensure a fair and reproducible evaluation, we employed a structured prompting strategy focusing on zero-shot prompting initially. Our methodology was designed to assess models’ inherent problem-solving abilities without additional external guidance [87]. Overall, for all evaluations, we incorporated a zero-shot prompting approach, where models were presented with mathematical problems without any prior exemplars or demonstrations. Zero-shot prompting is particularly valuable in assessing a model’s ability to generalize mathematical knowledge from its pretraining corpus and apply learned concepts to novel problems [87, 98]. By not providing explicit examples, this method evaluates how well a model can infer the appropriate solution methodology based solely on its prior learning.
Figure 2: Templates used to evaluate mathematical reasoning in LLMs. The multiple-choice format (a) ensures structured decision-making with predefined options, while the open-ended format (b) assesses free-form problem-solving capabilities.
[…]
For multiple-choice questions, we provided the set of answer choices and instructed the models to select the correct response from the given options, ensuring a structured decision-making process [68]. The prompt included clear formatting guidelines to standardize the response format across different models. In contrast, for open-ended problems that required free-form numeric or symbolic responses, we presented the question directly without predefined answer choices, allowing the models to generate responses based on their internal reasoning capabilities [81]. This approach ensured that multiple-choice tasks evaluated the models’ ability to differentiate between structured options, while open-ended problems tested their capability to derive solutions independently.
3.3.2 Sampling Settings
To ensure consistency and reliability in model responses, we employed a structured sampling strategy using temperature, top-k sampling, and top-p (nucleus) sampling [11, 70], which control randomness and diversity in token selection. To ensure a standardized evaluation across different LLMs, we utilized API access through OpenRouter for DeepSeek models and OpenAI API for GPT-4o, o1-mini, o1, and o3-mini, while maintaining their respective default generation settings. All models, including GPT-4o and o-series models, Gemini and DeepSeek use a default temperature of $T=1.0$ and an unrestricted probability sampling strategy (top-p = 1.0) [70], allowing the model to consider the full probability distribution for token selection.
Moreover, to ensure smooth execution of API requests while avoiding rate limits imposed by different providers, we implemented an adaptive time delay between successive API calls. This approach prevented throttling issues while maintaining efficiency in large-scale evaluation. All queries were formatted in a standardized manner to ensure fair comparisons across models.
3.3.3 Evaluation Metric
To assess the correctness of model-generated responses in mathematical tasks, we employ the Exact Match metric as our primary evaluation method [78]. Given that each question in our benchmark dataset has a single correct answer and the model produces a response per query, Exact Match ensures a rigorous evaluation by comparing the extracted answer to the ground truth.
[…]
Let $\hat{y}_{i}$ represent the extracted answer from the model’s output for the $i^{th}$ question, and let $y_{i}$ be the corresponding ground truth answer. The Exact Match accuracy is computed as:
[…]
where:
  • $N$ is the total number of evaluated questions.
  • $\mathbb{1}(\cdot)$ is the indicator function, returning $1$ if the extracted model response matches the ground truth after preprocessing, and 0 otherwise.
  • $\text{normalize}(\cdot)$ is a function that standardizes formatting, trims spaces, and normalizes numerical values.
Furthermore, error-handling mechanisms were incorporated to detect hallucinated responses [71, 79, 100], incomplete computations, or missing boxed outputs, with automated validation and manual review of flagged cases ensuring robust evaluation. Additionally, since some model responses exceeded token limits and resulted in truncated outputs, detection mechanisms were implemented to flag incomplete solutions, allowing either exclusion from evaluation or manual review when necessary.
Table 2: Comparison of ground truth and model-generated parsed answers
Reference answer Parsed Model answer Correctness
p - q [’p - q’] True
90$\circ$ [’90’] True
4 [’5’] False
[…]
6 - 5i [’6 - 5i’] True
Table 2 presents the evaluation of model-generated responses against ground truth values. The Reference Answer represents the ground truth, while the Parsed Model Answer contains the extracted response from the model’s output. The Correctness column indicates whether the model’s response exactly matches the ground truth after parsing, ensuring a strict evaluation of mathematical accuracy. Discrepancies, such as format mismatches or numerical errors, directly impact correctness scores, highlighting the model’s precision in solving mathematical problems.
By integrating automated parsing, formatting corrections, and evaluation metrics, our post-processing pipeline ensured reliable and unbiased assessment of mathematical reasoning capabilities across different LLM architectures.
4 Results
In this section, we conducted a quantitative and qualitative evaluation of the state-of-the-art LLMs mentioned earlier, using multiple public benchmark datasets for mathematical reasoning. As an initial assessment, we selected a subset of three widely recognized datasets [16, 39, 62], as discussed in Section 3, to systematically evaluate the performance of various cutting-edge models. Overall, our evaluation includes DeepSeek-R1 [56], DeepSeek’s most advanced reasoning model, along with its distilled variant DeepSeek-1.5B [37], which is derived from the R1 model. Additionally, we assessed Google’s latest Gemini 2.0 Flash [45] and four OpenAI models: GPT-4o, o1-mini, o1, and o3-mini [6, 44, 103]. Notably, while all these models are designed with reasoning capabilities, GPT-4o is the only one not explicitly optimized for providing reasoning-based CoT responses. These selected LLMs represent the most advanced AI systems available today, continuously evolving and competing across various mathematical reasoning benchmarks.
4.1 Evaluation of LLMs using MATH
As illustrated in Table 3, the performance of various LLMs on the MATH benchmark dataset [40] exhibits significant variability, underscoring differences in their mathematical reasoning capabilities.
In particular, GPT-4o [44] exhibited a significantly lower accuracy of 64.88%, lagging behind its peer models. This limitation may be attributed to deficiencies in complex reasoning capabilities or the inherent difficulty of competition-level mathematical problems in the dataset. In contrast, OpenAI’s o1 model achieved the highest accuracy (93.12%), closely followed by DeepSeek-R1 (90.45%), with both models surpassing the 90% correctness threshold—a benchmark that no other model reached. Other models demonstrated varying levels of performance: Gemini 2.0 Flash, 85.87%; o1-mini, 88.93%; o3-mini, 82.06%. However, a key observation is that the distilled variant of DeepSeek-R1 performed significantly worse, further reinforcing the hypothesis that model distillation can impair mathematical reasoning capabilities by compromising critical reasoning pathways in favor of improved computational efficiency.
Additionally, the quantitative comparisons across multiple LLMs, as presented in Figure 3, reveal consistent performance patterns across OpenAI’s models, while DeepSeek-R1 consistently outperformed its distilled variant. These findings highlight the trade-offs between model size and reasoning ability, as demonstrated by Fu et al. [33], underscoring the need for optimization techniques that enhance computational efficiency without sacrificing mathematical problem-solving proficiency.
Table 3: Performance comparison of various models on MATH in terms of accuracy.
Publisher Model Name Accuracy (%)
Google Gemini 2.0 Flash 85.87
OpenAI GPT-4o 64.88
o1-mini 88.93
o1 93.12
o3-mini 82.06
DeepSeek DeepSeek-1.5B 65.64
DeepSeek-R1 90.45
Figure 3: Performance comparison of various LLMs in the MATH benchmark. The accuracy of each model is displayed above the corresponding bar, highlighting differences in mathematical problem-solving capabilities across model publishers.
Furthermore, Figure 4 provides an in-depth comparison of the prompt responses and mathematical reasoning of DeepSeek-R1, o1, and Gemini 2.0 Flash. Interestingly, Gemini 2.0 Flash failed to provide a correct answer, whereas DeepSeek-R1 and o1 demonstrated more structured and logically coherent reasoning processes. This further illustrates the varying degrees of mathematical reasoning proficiency across models [103] and the importance of structured reasoning frameworks in achieving high accuracy.
Moreover, Figure 4 illustrates that DeepSeek-R1 and o1 followed nearly identical reasoning steps, correctly assigning coordinates, computing midpoints and centroids, and using determinant-based area calculations to arrive at the correct answer, 8. However, Gemini 2.0 Flash deviated from the correct reasoning path by making errors in its calculations. Specifically, despite correctly identifying some intermediate relationships, its handling of area scaling and altitude ratio calculations led to an incorrect final answer of 4 instead of, correct response of 8. This demonstrates a key limitation in Gemini 2.0 Flash’s numerical precision and reasoning accuracy for complex geometric problems compared with its peers.
Notably, the strong performance of DeepSeek-R1 in Figure 4, which relies heavily on RL methodologies [29], highlights the effectiveness of iterative, feedback-driven training paradigms in tackling mathematically intricate tasks. A crucial component in its training is Group Relative Policy Optimization (GRPO) [76], a more efficient RL approach than traditional methods. GRPO enables DeepSeek-R1 to excel in complex domains such as mathematics, science, and coding, reinforcing the potential of advanced learning techniques in enhancing model performance [37]. These findings underscore the importance of advanced reasoning mechanisms in solving competition-level mathematical problems, suggesting that RL frameworks-exemplified by DeepSeek-R1-represent a promising direction for enhancing the mathematical reasoning capabilities of next-generation LLMs.
Figure 4: Comparative evaluation of different LLMs on a Geometric problem from the MATH benchmark. The figure showcases step-by-step reasoning from three different models and highlights correctness. The question is shown in red, while the extracted answers are boxed. Green-highlighted sections indicate correct responses, whereas red-highlighted sections denote incorrect outputs. This comparison illustrates differences in mathematical reasoning and accuracy across models.
4.2 Evaluation of LLMs using GSM8K
As discussed in Section 3, the GSM8K benchmark dataset [16] comprises approximately 1.32k test samples designed to assess math reasoning in LLMs [16]. The zero-shot performance of various models is summarized in Table 4, presenting valuable insights into their mathematical reasoning capabilities.
In particular, the results revealed that DeepSeek-R1 [37] performed on par with OpenAI’s o1, with both models achieving the highest accuracy of 96.13%, surpassing all other peers in this evaluation. Additionally, several other models demonstrated strong performance: Gemini 2.0 Flash, 95.53%; GPT-4o, 95.58%; o1-mini, 95.91%; o3-mini, 95.83% However, a notable performance gap was observed in DeepSeek-R1’s distilled variant, which achieved only 81.13% accuracy—a significant decline compared to its full-scale counterpart. This performance disparity underscores the trade-offs associated with model distillation, where reducing the number of parameters for computational efficiency may inadvertently weaken logical reasoning and problem-solving capabilities. While distillation enables smaller, faster models, our findings suggest that excessive parameter reduction can compromise mathematical reasoning proficiency, potentially limiting their effectiveness in complex problem-solving tasks. Similarly, Figure 5 demonstrates that DeepSeek-R1 and o1 outperformed other peer LLMs. Notably, there are no significant differences across all LLMs released via OpenAI.
These results emphasize the delicate balance between model size and reasoning ability, highlighting the need for advancing distillation techniques that preserve core reasoning structures while maintaining computational efficiency.
Table 4: Performance comparison of various models on GSM8K in terms of accuracy.
Publisher Model Name Accuracy (%)
Google Gemini 2.0 Flash 95.53
OpenAI GPT-4o 95.98
o1-mini 95.91
o1 96.13
o3-mini 95.83
DeepSeek DeepSeek-1.5B 81.12
DeepSeek-R1 96.13
Moreover, the qualitative comparisons across multiple LLMs, as illustrated in Figure 6, reveal a noteworthy disparity in reasoning capabilities among OpenAI’s o1, DeepSeek-R1, and Google’s Gemini 2.0 Flash. The analysis demonstrates that DeepSeek-R1 exhibits superior performance in mathematical reasoning compared to its counterparts. Specifically, DeepSeek-R1 generated accurate solutions and also articulated a systematic, step-by-step process to arrive at its conclusions. In contrast, both OpenAI’s o1 and Gemini 2.0 Flash produced less comprehensive pathways, which resulted in erroneous answers. These findings, as shown in Figure 6, underscore DeepSeek-R1’s robust analytical proficiency relative to contemporary models, highlighting its advanced problem-solving capabilities in complex mathematical tasks.
Figure 5: Performance comparison of various LLMs in the GSM8K benchmark. The accuracy of each model is displayed above the corresponding bar, highlighting differences in mathematical problem-solving capabilities across model publishers.
Figure 6: Comparative evaluation of different LLMs on an Arithmetic problem from the GSM8K benchmark. The figure showcases step-by-step reasoning from three different models and highlights correctness. The question is shown in red, while the extracted answers are boxed. Green-highlighted sections indicate correct responses, whereas red-highlighted sections denote incorrect outputs. This comparison illustrates differences in mathematical reasoning and accuracy across models.
4.3 Evaluation of LLMs using MMLU
Leveraging the structured categorization of MMLU benchmark dataset [39], we conduct an in-depth comparative analysis of various LLMs across different mathematical domains.
First, as shown in Table 5, DeepSeek-R1 outperformed all peer models, achieving the highest accuracy of 97.62% in Formal Logic. Other models, such as o1, o3-mini, and Gemini 2.0 Flash, also demonstrated strong performance, all surpassing 90% accuracy. However, distilled variants of DeepSeek-R1 and o1-mini failed to reach this benchmark, reinforcing the hypothesis that model distillation can degrade logical reasoning capabilities by limiting access to essential computational pathways.
Table 5: Performance comparison of various models on MMLU in terms of accuracy.
Problem Type Model Name Accuracy (%)
Abstract Algebra Gemini 2.0 flash 87.00
GPT-4o 85.00
o1-mini 88.00
o1 94.00
o3-mini 96.00
DeepSeek-1.5B 61.00
DeepSeek-R1 94.00
College Mathematics Gemini 2.0 flash 96.00
GPT-4o 79.00
o1-mini 76.00
o1 99.00
o3-mini 99.00
DeepSeek-1.5B 65.00
DeepSeek-R1 96.00
Formal Logic Gemini 2.0 flash 90.48
GPT-4o 82.54
o1-mini 78.57
o1 95.24
o3-mini 96.03
DeepSeek-1.5B 47.62
DeepSeek-R1 97.62
High School Mathematics Gemini 2.0 flash 97.77
GPT-4o 88.15
o1-mini 98.89
o1 99.26
o3-mini 98.52
DeepSeek-1.5B 86.30
DeepSeek-R1 97.04
In addition, in Abstract Algebra, o3-mini outperformed its peers, achieving the highest accuracy of 96.00%. Abstract algebra is a particularly challenging subject, even for math graduates, as it demands a fundamental shift in mathematical thinking from traditional algebraic manipulations to analyzing complex algebraic structures such as groups, rings, and fields. Mastery of this subject requires a deep understanding of abstract concepts, rigorous proof construction, and advanced logical reasoning.
The results also suggest that OpenAI’s models, particularly o3-mini, excel in abstract reasoning and symbolic manipulation, enabling them to reconstruct and process complex algebraic structures effectively. In contrast, the distilled variant of DeepSeek-R1 performed significantly worse, achieving only 61% accuracy, further validating the detrimental impact of parameter reduction on logical inference and abstract reasoning.
Furthermore, in College Mathematics, both o1 and o3-mini achieved exceptional accuracy of 99.00% in College Mathematics, surpassing all other models. This outstanding performance can be attributed to two key factors: a). Superior reasoning capabilities: OpenAI’s o1 and o3-mini models are designed with enhanced mathematical reasoning frameworks, allowing them to efficiently process and solve complex college-level problems. b). Diversity of training data: The extensive dataset exposure during training may have equipped o1 and o3-mini with a broader mathematical foundation, leading to their unparalleled performance in advanced mathematical reasoning tasks. Meanwhile, Gemini 2.0 Flash also demonstrated strong performance, achieving a comparable accuracy of 96%, reinforcing its competence in college-level mathematics. However, distillation had a clear negative impact on performance, as evidenced by: 1). o1-mini, which achieved only 76.00% accuracy, a significant drop from its full-scale counterpart. 2). DeepSeek-1.5B, which attained just 65.00% accuracy, further emphasizing the trade-off between model size and reasoning ability.
This performance gap highlights the increased complexity of college mathematics compared to high school-level math. College-level coursework usually demands: 1). A deeper understanding of abstract mathematical concepts; 2). Stronger logical reasoning abilities; 3). Proficiency in complex proofs and real-world applications. Notably, the faster pace and heavier workload of college mathematics compared to high school math further contribute to its difficulty [7].
Moreover, given the relative simplicity of High School Mathematics, o1 achieved the highest accuracy at 99.62%, slightly improving upon its results in College Mathematics. Other peer models also performed well: DeepSeek-R1, 97.04%; o1-mini, 98.89%; o3-mini, 98.52%; Gemini 2.0 Flash, 97.77%. GPT-4o showed a lower accuracy (88.15%) compared with its peers in this domain too. Once again, the distilled variant of DeepSeek-R1 struggled, achieving only 86.30% accuracy, further demonstrating the performance degradation caused by model compression.
Similarly, Figure 7 presents a comprehensive comparative analysis of various LLMs across four key mathematical domains: Abstract Algebra, College Mathematics, Formal Logic, and High School Mathematics. Notably, OpenAI’s models demonstrate superior performance in abstract reasoning, particularly excelling in Abstract Algebra, where symbolic manipulation and deep theoretical understanding are crucial. This suggests that OpenAI’s models have strong capabilities in reconstructing and processing abstract mathematical structures.
Significantly, DeepSeek-R1 outperformed its peers in Formal Logic, showcasing its exceptional logical reasoning skills and ability to handle complex rule-based problem-solving more effectively than other models. These findings further highlight the specialized strengths of different LLMs and enhance the need for targeted model improvements to optimize performance across various mathematical domains.
Figure 7: Comparison of LLM models across different problem types in the MMLU benchmark dataset.
Figure 8: Comparative evaluation of different LLMs on a Formal Logic problem from the MMLU benchmark. The figure showcases step-by-step reasoning from three different models and highlights correctness. The question is shown in red, while the extracted answers are boxed. Green-highlighted sections indicate correct responses, whereas red-highlighted sections denote incorrect outputs. This comparison illustrates differences in mathematical reasoning and accuracy across models.
Figure 8 presents a comparative evaluation of different LLMs on a Formal Logic problem from the MMLU benchmark dataset [39]. The task involves identifying the correct conclusion from a given argument, with four answer choices labeled A, B, C, and D. The correct reference answer is option D, requiring models to distinguish between supporting premises and the actual conclusion. DeepSeek-R1 successfully identifies option D as the correct conclusion, offering a structured explanation by recognizing that the phrase "Because of this" serves as a linguistic connector rather than a part of the conclusion itself. In contrast, o1 and Gemini 2.0 Flash select option C, misinterpreting the logical structure of the argument and failing to differentiate between an explanatory transition and the central claim. The comparative analysis in Figure 8 further solidifies DeepSeek-R1’s exceptional logical reasoning skills and ability to handle complex rule-based problem-solving tasks.
Figure 9: Comparative evaluation of different LLMs on an Abstract Algebra problem from the MMLU benchmark. The figure showcases step-by-step reasoning from three different models and highlights correctness. The question is shown in red, while the extracted answers are boxed. Green-highlighted sections indicate correct responses, whereas red-highlighted sections denote incorrect outputs. This comparison illustrates differences in mathematical reasoning and accuracy across models.
Figure 9 illustrates a comparative evaluation of different LLMs in solving a problem related to abstract algebra, sourced from the MMLU benchmark dataset. The problem required reasoning about the number of elements of a given order in a group, specifically evaluating whether Statement 1 and Statement 2 hold true under group-theoretic principles. The task involved selecting the correct truth-value combination from four possible answer choices: A (True, True), B (False, False), C (True, False), and D (False, True). The reference answer, indicated in the Figure 9, is option A (True, True).
DeepSeek-R1 correctly identified both statements as true, providing a detailed justification using the properties of cyclic subgroups, Euler’s totient function, and disjoint subgroup structures. Similarly, o1 also arrived at the correct answer, correctly recognizing that exceeding 8 elements of order 15 necessitates at least 16 elements due to the existence of multiple distinct subgroups. However, Gemini 2.0 Flash incorrectly classified Statement 2 as false, demonstrating a reasoning flaw in its approach to subgroup structures and element counting. This comparison in Figure 9 highlights DeepSeek-R1 and o1’s superior reasoning abilities in abstract algebra, as both models successfully applied cyclic subgroup properties and disjoint set reasoning to reach the correct conclusion.
In summary, DeepSeek-R1 dominated in Formal Logic, while o3-mini excelled in Abstract Algebra. o1 and o3-mini outperformed all models in College Mathematics, likely due to their advanced reasoning frameworks and diverse training data. Model distillation significantly reduced performance as shown in [37], particularly in complex reasoning tasks like Abstract Algebra and College Mathematics. High school-level mathematics is relatively easier for LLMs, with multiple models surpassing 97% accuracy. These findings emphasize the trade-offs between model size, reasoning capability, and computational efficiency, highlighting the need for optimization strategies that maintain problem-solving proficiency while improving scalability.
4.4 Statistical Analyses
In this section, Table 6 presents a comparative evaluation of LLMs based on their mean accuracy and standard deviation across all three benchmark datasets. Among the evaluated models, OpenAI’s o1 achieves the highest mean accuracy (96.13%), followed closely by DeepSeek-R1 (95.21%) and o3-mini (94.57%). While Gemini 2.0 Flash attains a relatively high accuracy of 92.11%, it underperforms relative to the top-tier models.
In terms of stability, as quantified by standard deviation, both o1 (2.32) and DeepSeek-R1 (2.41) exhibit the most consistent performance. We believe that DeepSeek-R1’s high consistency can be attributed to its GRPO framework [37], which utilizes a multi-agent collaborative architecture, effectively integrates diverse reasoning pathways. This synergy helps mitigate task-specific biases and enhances consistency by reducing variability. The structured approach allows the model to dynamically prioritize high-confidence solutions while iteratively refining less certain outputs, thereby contributing to its robustness across heterogeneous datasets.
By contrast, o1-mini demonstrated significantly higher variability (standard deviation: 8.31), suggesting sensitivity to dataset-specific characteristics. Gemini 2.0 Flash (4.59) and o3-mini (5.74) exhibited moderate fluctuations in performance, likely due to limitations in their single-agent reasoning paradigms. GPT-4o also exhibited a significantly larger standard deviation (10.42) and was one of the two most inconsistent models alongside DeepSeek-1.5B. Meanwhile, DeepSeek-1.5B’s considerably lower mean accuracy (67.78%) and the highest standard deviation (12.82) reflect substantial variability in its responses, further cementing the detrimental impact of parameter reduction on mathematical reasoning.
Table 6: Performance statistics of different LLM models across all datasets, including mean accuracy (in %) and standard deviation.
LLM Models Gemini 2.0 Flash GPT-4o o1-mini o1 o3-mini DeepSeek-1.5B DeepSeek-R1
Mean (%) 92.11 82.59 87.72 96.13 94.57 67.78 95.21
Standard Deviation 4.59 10.42 8.31 2.32 5.74 12.82 2.41
Figure 10: Performance comparison of five LLMs across multiple datasets, grouped by publisher. The bars represent the mean accuracy (%), while the error bars indicate the standard deviation across three benchmark datasets.
Figure 10 presents the performance statistics of different LLMs across three datasets. The bar chart in Figure 10 illustrates the mean accuracy (%) of each model, with error bars representing the standard deviation of the accuracy percentages across datasets. The standard deviation bars indicate the variability in the performance of each model. Notably, o1 from OpenAI demonstrates the highest mean accuracy with a low variance, whereas DeepSeek-1.5B exhibits the largest variability, suggesting inconsistency in its accuracy across datasets. This visualization underscores that o1 and DeepSeek-R1 excel not only in accuracy but also in reliability for mathematical problem-solving tasks.
4.5 Response time analysis
Efficient response times are crucial in the deployment of LLMs, as they directly influence user experience and the practicality of real-time applications. Factors affecting inference speed include model size, computational complexity, hardware capabilities, and optimization techniques.
Figure 11: Latency Comparison across different LLMs
Notably, Figure 11 presents a comparative analysis of the average response time per query across various LLMs. The DeepSeek-R1 model exhibits the highest latency, averaging 81.0 seconds per response, which implies that DeepSeek-R1 is significantly slower than all other models. Among OpenAI’s models, o1 demonstrates a high response time of 15.0 seconds, likely due to its complex CoT process. GPT-4o also exhibits relatively high latency at 11.5 seconds, making it one of the three models that exceed the 10-seconds latency threshold. In contrast, o1-mini and o3-mini achieved faster inference speeds of 6.8 seconds and 5.3 seconds, respectively, outperforming the other two OpenAI models. DeepSeek-1.5B operates at 5.0 seconds, slightly outperforming OpenAI’s o3-mini models. Notably, Gemini 2.0 Flash exhibits the fastest response time at 4.2 seconds, making it the most efficient model in terms of inference speed.
Importantly, understanding these response times is essential for selecting appropriate models based on application requirements, balancing the trade-offs between latency and performance.
5 Discussion
In this study, we conducted a comprehensive evaluation of various LLMs to assess their mathematical reasoning capabilities across multiple benchmarks [16, 39, 40]. The evaluated LLMs include Gemini 2.0 Flash [45], GPT-4o [44], o1-mini [46], o1 [103], o3-mini [6], DeepSeek-1.5B [57], and DeepSeek-R1 [56]. Our empirical analysis revealed that DeepSeek-R1 and o1 outperformed its peers in most mathematical domains, demonstrating strong mathematical reasoning capabilities in solving structured problems. However, its performance was not on par with o3-mini in some complex and abstract fields, such as Abstract Algebra. Additionally, our findings underscore the impact of model distillation on mathematical reasoning performance. Specifically, distilled variants showcased noticeable declines in accuracy, reinforcing concerns that distillation, while improving computational efficiency, may inadvertently weaken the LLM’s capability to handle intricate problem-solving tasks. These observations suggest that there exists a trade-off between computational efficiency and reasoning depth, which needs to be carefully managed in future model optimizations.
First, our results further validate the advanced capabilities of DeepSeek-R1, which is trained using RL rather than Supervised Fine-Tuning (SFT) [31, 56, 69]. DeepSeek-R1 incorporates GRPO, a specialized RL-based optimization strategy that enhances training efficiency by evaluating model actions relative to a group of sampled peers. Unlike traditional RL approaches that require an external reward model, GRPO eliminates the need for a separate reward function by dynamically calculating advantages based on group-based scoring [56, 57]. This unique training paradigm delivers DeepSeek-R1 to balance cost and performance efficiently, leading to competitive accuracy in solving grade school-level mathematical problems. The superior performance of DeepSeek-R1 in GSM8K highlights its superiority in handling structured, multi-step arithmetic reasoning tasks. Notably, DeepSeek-R1 remains comparable to o1 and o3-mini, demonstrating that while GRPO enhances efficiency, it does not necessarily surpass models with specific CoT in all problem categories.
In addition, our analysis of DeepSeek-R1’s performance across MMLU benchmarks reveals a significant variation in mathematical reasoning capabilities across different domains. While DeepSeek-R1 excelled in Formal Logic, it struggled in more abstract mathematical areas, such as College Mathematics and Abstract Algebra. In particular, this discrepancy suggests that GRPO is particularly effective for solving problems that rely on structured logical reasoning, where a group-based approach to RL can refine logical consistency and decision-making [34]. Formal Logic problems, for instance, often involve pattern recognition and structured deduction [34], which could align well with GRPO’s framework. The group scoring mechanism in GRPO enables the model to internalize patterns commonly recognized by human reasoning, allowing it to perform well in problems that require rule-based logical deductions. However, the performance drop in Abstract Algebra and College Mathematics suggests that DeepSeek-R1 struggles with tasks requiring highly abstract conceptualization. Abstract Algebra, in particular, represents a fundamental shift from procedural algebra to the exploration of abstract mathematical structures such as groups, rings, and fields [99]. Unlike rule-based logic problems, abstract algebra necessitates deep symbolic manipulation, theorem proofs, and the ability to navigate complex algebraic structures [99], which require advanced deductive reasoning beyond GRPO’s core optimization strategy.
Furthermore, proof-based reasoning, a crucial component of higher-level mathematics, demands meta-cognitive processes that extend beyond pattern recognition, including recursive logic application and theorem proving. The limitations observed in DeepSeek-R1 to generalize across abstract mathematical problems indicate that a more versatile reasoning approach may be needed to complement GRPO-based training. In contrast, CoT reasoning demonstrates strong performance in specific mathematical fields by leveraging step-by-step logical breakdowns to enhance problem-solving accuracy. CoT excels in tasks that require explicit intermediate reasoning steps, which helps models navigate complex mathematical derivations more effectively. However, despite these advantages, CoT remains constrained by its tendency to overfit to structured problem-solving paradigms [9, 50], leading to limited adaptability when faced with problems requiring broader conceptual generalization. Thus, while CoT provides a deeper, structured reasoning framework, it may struggle to satisfy the generalization requirements necessary for handling diverse mathematical challenges. The trade-off between GRPO’s structured logic refinement and CoT’s sequential deductive approach suggests that a hybrid reasoning mechanism—integrating the strengths of both frameworks—may be essential for developing a more versatile and mathematically proficient LLM.
In summary, our findings align with prior research [2, 4, 27, 31, 47, 52, 54, 64, 66, 72, 76, 91, 103, 105, 107], further confirming that DeepSeek-R1 demonstrates strong mathematical reasoning compared to other LLMs. However, our research extends beyond previous work by providing a granular, field-specific analysis of LLM performance across different domains. Instead of evaluating general reasoning skills, we delve into the nuances of logical, arithmetic, and abstract reasoning capabilities, offering a holistic understanding of how different LLMs excel or struggle across mathematical disciplines. Despite its valuable contributions, our study has several limitations: 1). Limited Model Scope: While our evaluation includes prominent LLMs, we did not test other cutting-edge models, (such as Llama-3.2 / Llama-3.3, Claude, Mistral) which could provide additional insights into alternative training methodologies. 2). Restricted Benchmark Diversity: Our assessment relies on three primary benchmark datasets, which, while comprehensive, may not fully capture the breadth of mathematical reasoning challenges in real-world applications.
More importantly, future research should expand the scope of evaluated LLMs and incorporate additional benchmarks to gain a broader understanding of model strengths and weaknesses across varied mathematical disciplines. Specifically, integrating GRPO with peer CoTs could allow models to leverage both structured logical optimization and stepwise deductive reasoning. This integration could lead to LLM swarms, where multiple reasoning frameworks synergize to enhance efficiency and accuracy. Meanwhile, our study reaffirms that distillation can significantly impair reasoning capabilities. Future efforts should explore adaptive distillation methods that preserve deep reasoning pathways while optimizing for computational efficiency. Beyond current CoT frameworks [14, 53, 88, 89, 95], investigators could provide new insights into advanced reasoning frameworks, such as Chain of Draft (CoD) [93]. Those into how LLMs handle interconnected logical dependencies will improve abstract mathematical problem-solving [15]. Future models can strike a balance between efficiency, depth, and mathematical intelligence by advancing multi-framework LLMs and refining distillation strategies. Meanwhile, Xia et al. [88] introduced ReasonEval, a novel methodology for evaluating the quality of reasoning steps in mathematical tasks, emphasizing the need to assess both validity and redundancy in intermediate steps. Additionally, Frieder et al. [32] explored the mathematical capabilities of ChatGPT and GPT-4, demonstrating their effectiveness in querying mathematical facts and analyzing their performance across varying levels of mathematical complexity. These studies lay the foundation for more comprehensive and rigorous evaluations in the future, advancing our understanding of LLMs in mathematical reasoning.
6 Conclusion
Overall, our study presents a comprehensive comparative analysis of LLM performance in mathematical reasoning, identifying key strengths and weaknesses across multiple domains. We highlight the effectiveness of GRPO in structured problem-solving, demonstrating its ability to enhance logical consistency and rule-based reasoning. However, we also uncover its limitations in handling abstract mathematical concepts, where deep symbolic manipulation and theorem-based reasoning remain challenging for GRPO-trained models. Furthermore, our findings underscore the trade-offs associated with model distillation, revealing its potential drawbacks in preserving mathematical reasoning capabilities. As a result, we propose new strategies for refining distillation techniques, ensuring that compressed models maintain robust problem-solving proficiency without sacrificing computational efficiency. Beyond performance evaluation, this study paves the way for future advancements in LLM-driven mathematical intelligence by outlining potential enhancements in reasoning frameworks, including hybrid models that integrate RL with structured step-by-step inference methods. By addressing these challenges, our work contributes to the development of next-generation AI systems capable of handling complex mathematical reasoning with greater accuracy, efficiency, and adaptability.
7 Acknowledgement
The authors acknowledge the support of the Augusta University High Performance Computing Services (AUHPCS) for providing computational resources contributing to the results presented in this publication/report.

### v2

@@
-Unveiling the Mathematical Reasoning in DeepSeek Models: A Comparative Study of Large Language Models
+Evaluating Mathematical Reasoning Across Large Language Models: A Fine-Grained Approach
 Abstract
-With the rapid evolution of Artificial Intelligence (AI), Large Language Models (LLMs) have reshaped the frontiers of various fields, spanning healthcare, public health, engineering, science, agriculture, education, arts, humanities, and mathematical reasoning.
-Among these advancements, DeepSeek models have emerged as noteworthy contenders, demonstrating promising capabilities that set them apart from their peers.
-While previous studies have conducted comparative analyses of LLMs, few have delivered a comprehensive evaluation of mathematical reasoning across a broad spectrum of LLMs.
-In this work, we aim to bridge this gap by conducting an in-depth comparative study, focusing on the strengths and limitations of DeepSeek models in relation to their leading counterparts.
-In particular, our study systematically evaluates the mathematical reasoning performance of two DeepSeek models alongside five prominent LLMs across three independent benchmark datasets.
-The findings reveal several key insights: 1).
-DeepSeek-R1 consistently achieved the highest accuracy on two of the three datasets, demonstrating strong mathematical reasoning capabilities.
-2).
-The distilled variant of LLMs significantly underperformed compared to its peers, highlighting potential drawbacks in using distillation techniques.
-3).
-In terms of response time, Gemini 2.0 Flash demonstrated the fastest processing speed, outperforming other models in efficiency, which is a crucial factor for real-time applications.
-Beyond these quantitative assessments, we delve into how architecture, training, and optimization impact LLMs’ mathematical reasoning.
-Moreover, our study goes beyond mere performance comparison by identifying key areas for future advancements in LLM-driven mathematical reasoning.
-This research enhances our understanding of LLMs’ mathematical reasoning and lays the groundwork for future advancements.
+With the rapid advancement of Artificial Intelligence (AI), Large Language Models (LLMs) have significantly impacted a wide array of domains, including healthcare, engineering, science, education, and mathematical reasoning.
+Among these, mathematical reasoning remains a particularly challenging capability, often requiring multi-step logic and abstract generalization.
+While prior work has explored LLM performance on reasoning tasks, comprehensive evaluations that span both depth and breadth across model families remain limited.
+In this study, we present a systematic evaluation of mathematical reasoning abilities across eight leading LLMs, including two recent DeepSeek models, using three independent benchmark datasets.
+Our analyses reveal several key findings: (1) DeepSeek-R1 performs competitively with o1 across most domains and achieves the highest accuracy on the MMLU Formal Logic benchmark; (2) distilled variants, such as DeepSeek-1.5B, exhibit substantial performance degradation; and (3) Gemini 2.0 Flash achieves the lowest response latency.
+Beyond quantitative metrics, we explore how architectural choices, training paradigms, and optimization strategies contribute to variation in reasoning performance.
+These findings provide new insights into the capabilities and limitations of current LLMs in mathematical domains, and offer guidance for the development of future models better aligned with rigorous reasoning demands.
 1 Introduction
-As Artificial Intelligence (AI) continues to advance at an unprecedented pace, an array of powerful Large Language Models (LLMs) has emerged, including OpenAI’s GPT-4o, o1, o3, Claude 3.5, Llama 3.2, Qwen 2.5, and Gemini 2.0 [1, 12, 28, 30, 48, 57, 69, 82, 102].
-Driven by cutting-edge advancements in Deep Neural Networks (DNNs), these models integrate human cognition elements to enhance problem-solving and decision-making [49].
-Serving as a guiding light in natural language processing (NLP) [65], healthcare [21, 60, 61], clinical textual processing [22, 59, 104], biomedical image analyses [60], code generation [58], decision support [51], multimodal data analytics [83, 90], and mathematical reasoning [2, 52, 72, 76, 91, 103, 107], LLMs push the boundaries of AI capabilities, approximating human-like reasoning through sophisticated statistical inference [1, 48, 69].
+As Artificial Intelligence (AI) continues to advance at an unprecedented pace, an array of powerful Large Language Models (LLMs) has emerged, including OpenAI’s GPT-4o, o1, o3, Claude 3.5, Llama 4, Qwen 3, Gemini 2.0 and so on [Achiam et al., 2023, Chang et al., 2024, Dubey et al., 2024, Enis and Hopkins, 2024, Kasneci et al., 2023, Liu et al., 2024b, Team et al., 2024, Zhao et al., 2023].
+Driven by cutting-edge advancements in Deep Neural Networks (DNNs), these models integrate human cognition elements to enhance problem-solving and decision-making [Lappin, 2024].
+Serving as a guiding light in natural language processing (NLP) [Min et al., 2023], healthcare [Dai et al., 2023, Liu et al., 2023c, d], clinical textual processing [Dai et al., 2025, Liu et al., 2023b, Zhong et al., 2023], biomedical image analyses [Liu et al., 2023c], code generation [Liu et al., 2023a], decision support [Li et al., 2022], multimodal data analytics [Wang et al., 2024, Xiao et al., 2024], and mathematical reasoning [Ahn et al., 2024a, Li et al., 2024b, Peng et al., 2024, Shao et al., 2024, Xin et al., 2024, Zhong et al., 2024, Zhuang et al., 2024], LLMs push the boundaries of AI capabilities, approximating human-like reasoning through sophisticated statistical inference [Achiam et al., 2023, Kasneci et al., 2023, Neha and Bhati, 2025].
 However, despite their transformative potential, these models face notable limitations.
-Their high computational demands pose significant barriers to broader accessibility, making large-scale implementation costly [69].
+Their high computational demands pose significant barriers to broader accessibility, making large-scale implementation costly [Neha and Bhati, 2025].
 Moreover, while LLMs perform well in general contexts, they often struggle with specialized tasks, exhibiting inconsistencies in performance.
-Multimodal models, for example, continue to face challenges in spatial reasoning and real-world physics, while AI-assisted code generation frequently produces syntactically correct yet functionally flawed outputs, requiring human oversight [58, 103].
+Multimodal models, for example, continue to face challenges in spatial reasoning and real-world physics, while AI-assisted code generation frequently produces syntactically correct yet functionally flawed outputs, requiring human oversight [Liu et al., 2023a, Zhong et al., 2024].
 These constraints underscore the ongoing need for refinement and innovation in AI research to bridge the gap between artificial and human intelligence.
-In particular, GPT-4o, released by OpenAI in May 2024, is a multimodal model capable of processing text, images, and voice with remarkable efficiency.
-Leveraging an advanced transformer architecture, it surpasses GPT-3 in critical areas such as mathematical reasoning and language comprehension [1, 23, 44, 75].
-With an estimated 2 trillion parameters, GPT-4o is significantly larger than its predecessors, enabling substantial improvements in performance and adaptability.
-Meanwhile, other cutting-edge models, such as o1 and o3, have been introduced in 2024 and 2025, respectively.
-In particular, o1 enhances reasoning capabilities by incorporating a Chain-of-Thought (CoT) approach [80], where it renders intermediate reasoning steps before reaching a final answer closely mirroring the way humans process complex problems [103].
-More significantly, o3 advances reasoning performance through a "simulated reasoning" process, in which the model actively generates and evaluates multiple solution paths [6].
-By pausing, reflecting, and adjusting its approach before delivering a final answer, o3 exhibits human-like reasoning, allowing for more nuanced and adaptive problem-solving, particularly in highly complex scenarios [67].
-However, none of the GPT models, including GPT-4o, o1, and o3, are open-source, as they remain proprietary.
-In addition, Claude 3.5, released in 2024, is built on previous versions.
-It emphasizes safety, alignment, and performance, with improvements in reasoning, language understanding, and handling complex tasks like text and code generation [30].
+In 2024, OpenAI released GPT-4o, a multimodal model capable of processing text, images, and audio.
+With estimated 2 trillion parameters, it significantly surpasses GPT-3 in tasks such as reasoning, coding, and clinical diagnostics [Achiam et al., 2023, Dale, 2021, Hurst et al., 2024, Shahriar et al., 2024].
+Meanwhile, other cutting-edge models, such as o1, o3 and recently o4-mini, have been introduced between 2024 and 2025.
+In particular, o1 enhances reasoning capabilities by incorporating a Chain-of-Thought (CoT) approach [Stechly et al., 2025, Zhong et al., 2024], where it renders intermediate reasoning steps before reaching a final answer closely mirroring the way humans process complex problems [Arrieta et al., 2025].
+In addition, Claude 3.5, released in 2024, emphasizes safety, alignment, and performance, with improvements in reasoning, language understanding, and handling complex tasks like text and code generation [Enis and Hopkins, 2024].
 With 250B parameters, it surpasses earlier models in accuracy and ethical alignment.
-It supports up to 200K tokens for extended context, enabling better processing of larger inputs.
-Notably, enhanced by reinforcement learning from human feedback (RLHF) [41, 85] and Constitutional AI, it reduces undesirable responses, biases, and better aligns with human intent.
-Claude 3.5 excels in specialized areas like coding and scientific reasoning, with improved transparency and ethical safeguards [30, 69].
+It supports up to 200K tokens for extended context.
+Notably, enhanced by reinforcement learning from human feedback (RLHF) [Hu et al., 2024, Wen et al., 2024] and Constitutional AI, it reduces undesirable responses, biases, and better aligns with human intent.
+Claude 3.5 excels in specialized areas like coding and scientific reasoning, with improved transparency and ethical safeguards [Enis and Hopkins, 2024, Neha and Bhati, 2025].
 Furthermore, Llama-3.3, presented in 2024, is the latest version of LLM in Meta AI family, following Llama-1 and Llama-2.
-Llama-3.3 advances further with 70B parameters and 128K token context window, improved by grouped-query attention for better efficiency [28, 63].
+Llama-3.3 advances further with 70B parameters and 128K token context window, improved by grouped-query attention for better efficiency [Dubey et al., 2024, Lu et al., 2024].
 Llama 3.1 excels in coding, logical problem solving, and low-resource language tasks.
-Unlike closed models such as GPT series, it remains open-weight and freely accessible for research and commercial use, but is restricted to text-only input [26, 63, 69].
-Safety measures, such as automated red-teaming and filtered training data, help minimize undesirable outputs.
-Moreover, Gemini 2.0 Flash is the latest multimodal LLM of Google, building on versions 1.0 and 1.5 to offer more robust generative AI capabilities across text, images, audio, and video.
-Moreover, Gemini 2.0 Flash, initially introduced as an experimental variant, provides significant speed and efficiency gains over its predecessor, Gemini 1.5 Flash, without sacrificing the efficiency [43, 45, 82].
+Unlike closed models such as GPT series, it remains open-weight and freely accessible for research and commercial use, but is restricted to text-only input [Dominguez-Olmedo et al., 2025, Lu et al., 2024, Neha and Bhati, 2025].
+Moreover, Gemini 2.0 Flash, initially introduced as an experimental variant, provides significant speed and efficiency gains over its predecessor, Gemini 1.5 Flash, without sacrificing the efficiency [Huang and Zhang, 2024, Imran and Almusharraf, 2024, Team et al., 2024].
 Notably, it outperforms Gemini 1.5 Pro on key benchmarks while operating at twice the speed.
@@
 By combining better performance in tasks such as math, code generation, and multilingual audio output with enhanced efficiency.
-Gemini 2.0 aims to deliver comprehensive, cost-effective AI solutions for both developers and end users [43, 82].
+Gemini 2.0 aims to deliver comprehensive, cost-effective AI solutions for both developers and end users [Huang and Zhang, 2024, Team et al., 2024].
 Lastly, Qwen2.5, released in September 2024, is the latest iteration in the Qwen series, following Qwen2 in June 2024 and the original Qwen in August 2023.
 Qwen1.5 featured models up to 72B parameters, emphasizing efficiency and open-source accessibility.
-Qwen2 introduced improved reasoning, multilingual support, and coding capabilities, with models scaling up to 72.71B parameters.
-[94, 106]
-Besides, established in 2023 as a research initiative to push the boundaries of artificial general intelligence (AGI), DeepSeek models set out to overcome existing limitations by developing specialized models focused on efficiency, adaptability, and domain expertise [8, 20, 38, 37, 57, 48, 56, 76].
-In 2024, the Mixture-of-Experts (MoE), an efficiency-driven architecture that leverages sparse activation to reduce computational overhead [38, 56, 57], for DeepSeek was introduced.
-This was followed by the launch of DeepSeek Coder, a suite of code-focused models ranging from 1B to 33B parameters, designed to streamline software development workflows.
-Meanwhile, DeepSeek Math, trained on 120B math-related tokens, was developed to handle advanced mathematical and symbolic reasoning tasks [76].
-Expanding its model portfolio, DeepSeek introduced the V2 and V3 series [56, 57].
-V2 implemented multi-head latent attention (MLA) alongside a MoE system with 236B parameters in total, of which only 21B were active per query, optimizing computational efficiency [56].
-V3, an open-source model, further enhanced efficiency with 671B total parameters, activating only 37B per query, excelling in complex reasoning tasks while minimizing resource demands and reliance on supervised data [57].
+Qwen2 introduced improved reasoning, multilingual support, and coding capabilities, with models scaling up to 72.71B parameters [Yang et al., 2024, Zhu et al., 2025].
+Besides, established in 2023 as a research initiative to push the boundaries of artificial general intelligence (AGI), DeepSeek models set out to overcome existing limitations by developing specialized models focused on efficiency, adaptability, and domain expertise [Bi et al., 2024, Dai et al., 2024, Guo et al., 2024, 2025, Liu et al., 2024b, Kasneci et al., 2023, Liu et al., 2024a, Shao et al., 2024].
+In 2024, the Mixture-of-Experts (MoE), an efficiency-driven architecture that leverages sparse activation to reduce computational overhead was introduced [Guo et al., 2024, Liu et al., 2024a, b].
+Meanwhile, DeepSeek Math, trained on 120B math-related tokens, was developed to handle advanced mathematical and symbolic reasoning tasks [Shao et al., 2024].
 In 2025, DeepSeek made a significant breakthrough with R1 Zero, incorporating self-verification, reflection, and extended CoTs.
-Notably, DeepSeek-R1 was recently presented, specifically designed for mathematical, coding, and logical problem-solving, enhancing autonomous decision-making and precision in both research and enterprise applications [37, 38, 76].
-To extend accessibility, an open-sourced suite of distilled models for DeepSeek, optimized for deployment in resource-constrained environments such as edge computing platforms and low-memory systems.
-These models preserve scalability and cost-effectiveness, making cutting-edge AI more accessible across diverse applications.
-Notably, mathematical reasoning usually poses intricate challenges without straightforward solutions [3, 55].
+Notably, DeepSeek-R1 was recently presented, specifically designed for mathematical, coding, and logical problem-solving, enhancing autonomous decision-making and precision in both research and enterprise applications [Guo et al., 2025, 2024, Shao et al., 2024].
+Notably, mathematical reasoning poses intricate challenges without straightforward solutions [Ahn et al., 2024b, Liang et al., 2024].
 Unlike routine tasks that follow established frameworks, mathematical reasoning demands creativity, abstract thinking, and advanced cognitive skills.
-In this study, we systematically evaluate the mathematical reasoning performance of a wide array of LLMs, with a particular emphasis on DeepSeek models, to further unveil machine creativity in mathematical reasoning.
+In this work, we systematically evaluate the mathematical reasoning performance of a wide array of LLMs, with a particular emphasis on DeepSeek models, to further unveil machine creativity in mathematical reasoning.
 Leveraging multiple independent datasets and quantitative metrics, we conduct a comprehensive empirical analysis to assess and compare their capabilities.
-Our findings provide a detailed evaluation of prominent LLMs in mathematical reasoning while highlighting the advancements and unique strengths of DeepSeek models.
-2 Related Works
-Recent research has increasingly focused on summarizing and evaluating the performance of LLMs in problem-solving and code generation, areas where general-purpose LLMs excel in text-based tasks but often struggle with mathematical precision and structured reasoning [24, 36, 48, 77, 86, 97].
-To address these limitations, LLMs have prioritized enhancing reasoning capabilities and improving computational efficiency in next-generation models, aiming to bridge the gap between linguistic fluency and robust problem-solving skills.
-Specifically, in 2024, Wang et al., summarized and detailed the specifications of Multimodal Large Language Models (MLLMs), standing out at the forefront of AI [83].
-In brief, Wang et al., overviewed a wide array of data modalities, such as text, images, audio, and sequential data [83].
-In fact, MLLMs play a crucial role in multimodal understanding tasks by integrating text and image information to achieve more intelligent and comprehensive understanding and reasoning.
-In this field, Wang et al., outlined the developments, advancements, and utilization of multiple MLLMs, including MiniGPT-4, InstructBLIP, and Wiki-LLaVA are three highly regarded models, along with other related MLLMs such as 3DMIT, GroundingGPT, ModaVerse, Vary-toy, LLaVAMOLE, and CogCom [83].
-In addition, at the end of 2024, Zhong et al.
-established a comprehensive and extensive evaluation of o1-preview (the early version of o1) to showcase the advancement of OpenAI’s o1-preview across a diverse array of complex reasoning tasks, spanning multiple domains, including computer science, mathematics, natural sciences, medicine, linguistics, and social sciences [103].
-Through extensive testing, o1-preview highlighted remarkable capabilities, often achieving human-level or superior performance in areas ranging from coding challenges to scientific reasoning and from language processing to creative problem-solving [103].
-Key findings include: 1).
-83.3% success rate in solving complex competitive programming problems, surpassing human experts.
-2).
-Superior ability in generating coherent and accurate radiology reports, outperforming other evaluated models.
-3).
-100% accuracy in high school-level mathematical reasoning tasks, providing detailed step-by-step solutions.
-4).
-Advanced natural language inference capabilities across general and specialized domains like medicine.
-5).
-Impressive performance in chip design tasks, outperforming specialized models in areas such as script generation and bug analysis.
-6).
-Remarkable proficiency in anthropology and geology, demonstrating deep understanding and reasoning in these specialized fields.
-7).
-Comprehensive financial knowledge and strong capabilities in quantitative investing.
-8).
-Effective performance in social media analysis, including sentiment analysis and emotion recognition.
-9).
-Excellent performance in educational measurement and psychometrics, demonstrating a solid grasp of standard psychometric concepts that are equivalent to or beyond a first-year master’s or doctoral student’s level.
-Furthermore, Yang et al.
-investigated the capabilities of advanced LLMs, particularly the o1 model, in literary analysis [96].
-Given the prestige of the Nobel Prize and its emphasis on cultural, historical, and linguistic depth, applying LLMs to these works offers valuable insights into both human and AI approaches to literary interpretation [96].
-The study employed qualitative and quantitative evaluations to assess coherence, creativity, and fidelity to the text, shedding light on the strengths and limitations of AI in domains traditionally dominated by human expertise.
-While LLMs demonstrated remarkable analytical capabilities, particularly in structured tasks, they struggled with emotional nuance and coherence—areas where human interpretation remains unparalleled.
-This research underscores the transformative potential of human-AI collaboration in the humanities, paving the way for new opportunities in literary studies, textual analysis, and interdisciplinary research.
-By leveraging AI’s analytical power alongside human intuition, this study highlights promising directions for AI-assisted literary interpretation and beyond.
-Moreover, Xu et al.
-demonstrated that LLMs have made significant advancements in clinical decision-making, particularly those leveraging in-context demonstrations and specialized medical fine-tuning [92].
-These models exhibit strong performance in medical language processing but continue to face challenges in real-time adaptability, multi-step reasoning, and handling complex medical tasks.
-Agent-based AI systems aim to overcome these limitations by incorporating reasoning traces, contextual tool selection, knowledge retrieval, and both short- and long-term memory.
-These features enable medical AI agents to manage complex clinical scenarios where decision-making requires real-time interaction with the environment [92].
-Unlike conventional model-based approaches that treat medical queries as isolated questions, medical AI agents approach them as dynamic, multi-faceted tasks, allowing them to function more like human doctors.
-Xu also investigated the selection of the backbone LLM for medical AI agents, which serves as the foundation for their reasoning and action generation.
-Specifically, Xu et al.
-[92] examined the capabilities of the emerging o1 model and its impact on agents’ reasoning, tool-use adaptability, and real-time information retrieval across diverse clinical scenarios, including high-stakes environments such as intensive care units (ICUs).
-The findings highlighted o1’s potential to enhance diagnostic accuracy and consistency, paving the way for more intelligent, responsive AI tools that support improved patient outcomes and more effective decision-making in clinical practice [92].
-Notably, Ahn et al.
-[3] addressed four critical aspects of LLMs in mathematical reasoning: 1).
-mathematical problem types and datasets, 2).
-techniques for enhancing LLM performance (such as prompt engineering and fine-tuning), 3).
-factors affecting model effectiveness (including scale and pre-training data), and 4).
-challenges such as brittleness and inconsistency, which often cause models to generate different answers for similar problems.
-Besides, Liang et al.
-presented a comprehensive overview on the applications of AI in mathematical research, highlighting the transformative role AI has begun to play in this domain [55].
-Traditionally, AI advancements have heavily relied on theoretical foundations provided by mathematics and statistics.
-However, recent developments in AI, particularly in reinforcement learning (RL) and LLMs, have demonstrated the potential for AI to contribute back to mathematics by offering flexible algorithmic frameworks and powerful inductive reasoning capabilities that support various aspects of mathematical research.
-This survey aimed to establish a bridge between AI and mathematics, providing insights into the mutual benefits and fostering deeper interdisciplinary understanding.
-In particular, Liang et al.
-discussed that while current AI and LLMs may struggle with complex deductive reasoning, their inherent creativity, the capability to generate outputs at high throughput based on recognition of shallow patterns, holds significant potential to support and inspire mathematical research [55].
-Furthermore, Liang et al.
-[55] addressed the lack of cross-disciplinary communication: mathematicians may not fully comprehend the latest advances in AI, while AI researchers frequently prioritize benchmark performance over real-world applications in frontier mathematical research.
-Lastly, this work seeks to close that gap, offering a detailed exploration of AI fundamentals, its strengths, and its emerging applications in the mathematical sciences.
-While various studies have begun exploring in-depth comparisons between DeepSeek-R1 and other leading models [37, 38, 76], few provide a comprehensive analysis of their mathematical reasoning capabilities.
-Thus, in this work, we introduce a holistic empirical study to systematically compare the mathematical reasoning performance of DeepSeek-R1 with its peer models.
-Our goal is to contribute to a deeper understanding of DeepSeek-R1’s strengths and limitations while inspiring further advancements in the future works, using a wide array of peer models, datasets, and quantitative metrics.
-3 Methodology
-In this section, we provide a comprehensive overview of our experimental framework, detailing the datasets utilized, the quantitative evaluation metrics, as well as the overall methodology.
+Our findings render a detailed evaluation of prominent LLMs in mathematical reasoning while highlighting the advancements and unique strengths of LLMs.
+2 Experimental Framework
+In this section, we provide a comprehensive evaluation of our experimental framework, detailing the datasets utilized, the quantitative evaluation metrics, as well as the overall methodology.
 We also present a justification for selecting these datasets as benchmarks, emphasizing their relevance, diversity, and suitability in effectively assessing LLM performance in mathematical reasoning.
 By establishing a consistent evaluation framework, we aim to ensure a thorough and insightful comparison of model capabilities across various reasoning tasks.
-Figure 1: This figure illustrates the experimental framework of this work, including five vital components.
-3.1 Experimental Frameworks
-The framework of this empirical study is illustrated in Figure 1 and consists of five key components: 1).
-Benchmark Data: We utilized three representative benchmark datasets [16, 39, 62] to evaluate the mathematical reasoning capabilities of multiple LLMs.
-2).
-Zero-Shot Prompting: Zero-shot prompting was employed to design all prompts [42], ensuring a standardized evaluation approach.
-3).
-Language Model Selection: A diverse set of representative LLMs was selected, including Google’s Gemini 2.0 Flash [45], OpenAI’s GPT-4o [1], o1-mini [46], o1 [103], o3-mini [6], DeepSeek’s DeepSeek-R1-Distill-Qwen-1.5B [37], hereafter referred to as DeepSeek-1.5B, and DeepSeek-R1 [56].
-4).
-Evaluation: Accuracy was used as a critical quantitative metric to assess and compare the mathematical reasoning performance of these models [17, 18, 19].
-5).
-Comparison Between Models: A comprehensive statistical analysis was conducted to evaluate and contrast the performance of each LLM, offering deeper insights into their mathematical reasoning capabilities.
-Overall, this structured framework enabled a holistic and systematic comparison of LLMs in mathematical reasoning.
+Figure 1: Overview of the experimental framework: (A) Task preparation involves selecting benchmark datasets spanning academic, graduate-level, and competition math problems.
+(B) Zero-shot prompting is used to query various LLMs with structured instructions.
+(C) Model outputs are evaluated using DeepEval across reasoning, clarity, and correctness dimensions, enabling quantitative comparison of model performance.
+2.1 Experimental Designs
+The framework of this empirical study is illustrated in Figure 1 and consists of three sections: A) We selected benchmarks consisting of diverse math datasets.
+[Cobbe et al., 2021, Hendrycks et al., 2020, Lu et al., 2023] B).
+Zero-Shot Prompting: Zero-shot prompting was employed to design all prompts [Hu et al., 2025], ensuring a standardized evaluation approach.
+The prompts are uniformly applied across a suite of language models.
+[Imran and Almusharraf, 2024, Achiam et al., 2023, Guo et al., 2025, Arrieta et al., 2025, Zhong et al., 2024, Liu et al., 2024a](C) We evaluate model outputs using DeepEval’s [Ip and Vongthongsri, 2025] fine-grained metrics: correctness, clarity, and reasoning, each computed by comparing predicted answers to ground truth under structured evaluation steps.
+Overall, this structured framework enabled holistic and systematic evaluations of LLMs in mathematical reasoning.
 2.2 Datasets
 To evaluate the mathematical reasoning capabilities of LLMs, we employ well-established benchmark datasets that encompass diverse mathematical domains, question types, and difficulty levels.
-The selected datasets include Math Competition (MATH) [40], Grade School Math 8K (GSM8K)  [16], and Massive Multitask Language Understanding (MMLU) [39], covering topics ranging from elementary arithmetic to advanced algebra, logic, and mathematical competitions.
+The selected datasets include Math Competition (MATH) [Hendrycks et al., 2021], Grade School Math (GSM8K) [Cobbe et al., 2021], and Massive Multitask Language Understanding (MMLU) [Hendrycks et al., 2020] math subset, covering topics ranging from elementary arithmetic to advanced algebra, logic, and mathematical competitions.
 From each benchmark, we carefully select a representative subset of problems to ensure a balanced evaluation while maintaining the diversity and complexity of mathematical reasoning tasks.
-These subsets are chosen for their broad coverage of mathematical domains, variety in question formats, and representation of different complexity levels, ensuring a comprehensive assessment of LLMs’ symbolic manipulation, logical reasoning, and problem-solving abilities [35, 74, 101].
-In total, our evaluation spans 2,178 mathematical problems across multiple benchmark datasets, ensuring a comprehensive assessment of LLMs’ mathematical reasoning capabilities.
-By evaluating LLMs on these benchmarks, we identified their strengths and limitations in mathematical reasoning, aiding in the development of more robust and capable models.
-Table 1 provides an overview of the datasets utilized in this study.
+These subsets are chosen for their broad coverage of mathematical domains, variety in question formats, and representation of different complexity levels, ensuring a comprehensive assessment of LLMs’ symbolic manipulation, logical reasoning, and problem-solving abilities [Gema et al., 2024, Setlur et al., 2024, Zhang et al., 2024].
+Figure 2: Data distribution across selected datasets for evaluating mathematical reasoning.
+Notably, these subsets are chosen for their broad coverage of mathematical domains, variety in question formats, and representation of different complexity levels, ensuring a comprehensive assessment of LLMs’ logical reasoning, and problem-solving capabilities [Gema et al., 2024, Setlur et al., 2024, Zhang et al., 2024].
+We selected 262 free-form numeric problems from MATH dataset (mathematics competitions), covering domains such as algebra, geometry, number theory, and combinatorics.
+For this study, we selected problems spanning $\geq 3$ difficulty levels to ensure coverage between medium and hard tasks.
+The GSM8K testset includes 1,320 arithmetic word problems designed for the grade school level.
+From MMLU, we incorporate four subject-specific subsets—College Mathematics (MCM), Abstract Algebra (MAA), Formal Logic (MFL), and High School Mathematics (MHM)—ranging from 100 to 270 multiple-choice questions.
+These subsets cover diverse topics such as calculus, group theory, logical reasoning, and general high school math.
+In total, our experimental benchmark includes 2,178 problems, offering broad coverage across difficulty levels, formats, and mathematical domains to support fine-grained evaluation of reasoning skills in LLMs.
+By evaluating LLMs on these benchmarks, we identified their strengths and limitations in mathematical reasoning, aiding in the development of more robust and capable models Figure 2 presents an overview structure of the datasets utilized in this study.
+The detailed table of the dataset is shown in Appendix A.
+2.3 Evaluation of Mathematical Reasoning Across LLMs
+As discussed previously, multiple LLMs have demonstrated remarkable capabilities in mathematical reasoning [Zhou et al., 2024].
+To systematically and comprehensively evaluate the mathematical reasoning capabilities of various LLMs, we conducted experiments using API-based access to a diverse set of models.
+The selected models include OpenAI’s GPT-4o, o1-mini, o1, and o3-mini, DeepSeek’s DeepSeek-1.5B and DeepSeek-R1, Google’s Gemini 2.0 Flash, and Meta’s Llama-4 the latest one.
+These models were accessed through their respective API services using secure API keys from Google, OpenAI, DeepSeek, and Openrouter.
+2.3.1 Prompting Strategy
+The effectiveness of LLMs in mathematical reasoning is highly dependent on the design of input prompts [Amatriain, 2024, Cain, 2024, Denny et al., 2024].
+To ensure a fair and reproducible evaluation, we employed a structured prompting strategy focusing on zero-shot prompting initially.
+Our methodology was designed to assess models’ inherent problem-solving abilities without additional external guidance [Wu et al., 2024].
+Overall, for all evaluations, we incorporated a zero-shot prompting approach, where models were presented with mathematical problems without any prior exemplars or demonstrations.
+Zero-shot prompting is particularly valuable in assessing a model’s ability to generalize mathematical knowledge from its pretraining corpus and apply learned concepts to novel problems [Wu et al., 2024, Yuan et al., 2024].
+By not providing explicit examples, this method evaluates how well a model can infer the appropriate solution methodology based solely on its prior learning.
+For multiple-choice questions, we provided the set of answer choices and instructed the models to select the correct response from the given options, ensuring a structured decision-making process [Myrzakhan et al., 2024].
+In contrast, for open-ended problems that required free-form numeric or symbolic responses, we presented the question directly without predefined answer choices, allowing the models to generate responses based on their internal reasoning capabilities [Tam et al., 2025].
+This approach ensured that multiple-choice tasks evaluated the models’ ability to differentiate between structured options.
+The prompt templates linked with each category question are elaborated in Appendix B.
+2.3.2 Settings
+To ensure a standardized evaluation across different LLMs, we utilized API access through OpenAI and OpenRouter, while maintaining their respective default generation settings.
+All models, including GPT-4o and o-series models, Gemini, Llama-4 and DeepSeek use a default temperature of $T=1.0$ and an unrestricted probability sampling strategy (top-p = 1.0) [Nguyen et al., 2024], allowing the model to consider the full probability distribution for token selection.
+2.3.3 Evaluation Metric
+We evaluate the model outputs using three core metrics, Correctness, Clarity, and Reasoning, which reflect different dimensions of mathematical understanding and communication.
+The evaluation is conducted using LLM-as-a-judge methodology implemented via the DeepEval framework.[Ip and Vongthongsri, 2025]
+Correctness measures whether the model output reaches the right final answer or conveys the correct mathematical content, allowing for variations in formatting.
+Clarity assesses the logical coherence and readability of the explanation, ensuring that the reasoning process is easy to follow.
+Reasoning evaluates the depth and validity of the model’s step-by-step problem-solving approach.
+Each model-generated response is processed through a pre-defined prompt template with access to the original question and reference answer.
+DeepEval internally prompts a trusted evaluator model (e.g., GPT-4o) to return scalar ratings for each metric.
+Let $M$ be a model, $Q$ be a set of questions, and $\text{Score}_{M}^{(k)}(q)$ be the score for metric $k\in\{\text{Correctness},\text{Clarity},\text{Reasoning}\}$ on question $q\in Q$.
+The average metric score is computed as:
+$\text{Score}_{M}^{(k)}=\frac{1}{|Q|}\sum_{q\in Q}\text{LLM}_{\text{eval}}\left(\text{prompt}(q,\text{output}_{M}(q),\text{gold}(q),\text{steps}_{k})\right)$
+3 Results and Analysis
+In this section, we conducted a quantitative and qualitative evaluation of the state-of-the-art LLMs mentioned earlier for mathematical reasoning.
+As an initial assessment, we selected a subset of three widely recognized datasets [Cobbe et al., 2021, Hendrycks et al., 2020, Lu et al., 2023], as introduced in Section 3, to systematically evaluate the performance of various cutting-edge models.
+Overall, our evaluation in Table 3, covering diverse evaluation metric, such as Accuracy(Acc), Reasoning Score(RS), and Clarity Score(CS), includes DeepSeek-R1 [Liu et al., 2024a],along with its distilled variant DeepSeek-1.5B [Guo et al., 2025], which is derived from the R1 model.
+Additionally, we assessed Google’s latest Gemini 2.0 Flash [Imran and Almusharraf, 2024], four OpenAI models: GPT-4o, o1-mini, o1, and o3-mini [Arrieta et al., 2025, Hurst et al., 2024, Zhong et al., 2024] and finally Meta’s Llama-4 the latest model [AI, 2025].
+Notably, while all these models are designed with reasoning capabilities, GPT-4o and Llama-4 are not explicitly optimized for providing reasoning-based CoT responses but they can still perform CoT reasoning quite well when prompted, especially in instruction-tuned variants.
+These selected LLMs represent the most advanced AI systems available today, continuously evolving and competing across various mathematical reasoning benchmarks.
+Table 1: Performance comparison of LLMs across multiple mathematical reasoning abilities using three metrics: Accuracy (ACC), Reasoning Score (RS), and Clarity Score (CS).
+The results highlight the average performance of each model across all categories.
+LLMs Math GSM8K MAA MFL MCM MHM Avg.
+[…]
+[…]
+Deepseek-R1 92.37 96.36 95.00 97.62 96.00 97.78 95.75
+Deepseek-1.5B 68.70 81.58 61.00 52.38 66.00 88.15 67.25
+o1 95.80 96.21 94.00 96.51 99.00 99.26 96.91
+o1-mini 89.69 96.13 88.00 78.57 89.29 98.89 88.89
+o3-mini 96.56 94.84 96.00 95.24 99.00 98.52 97.06
+Gemini-2.0-Flash 87.02 94.24 88.00 90.48 96.00 97.77 91.85
+Llama-4 81.68 97.04 88.00 90.48 93.00 97.41 90.11
+[…]
+[…]
+Deepseek-R1 95.15 98.15 94.76 94.04 98.26 98.19 96.08
+Deepseek-1.5B 77.58 88.56 72.54 66.95 80.82 89.33 77.64
+o1 98.73 98.82 94.00 86.84 99.09 98.20 95.97
+o1-mini 95.51 98.46 87.49 95.63 93.09 98.51 94.04
+o3-mini 99.04 97.53 97.37 98.44 99.05 98.33 98.45
+Gemini-2.0-Flash 89.52 96.01 93.54 93.78 92.07 97.12 93.60
+Llama-4 92.34 98.15 93.52 95.59 95.54 97.99 94.59
+[…]
+[…]
+Deepseek-R1 85.89 97.21 87.81 89.13 84.42 88.10 87.07
+Deepseek-1.5B 77.21 88.33 60.09 66.27 80.27 83.27 73.62
+o1 86.80 98.34 84.91 77.76 85.00 87.95 84.48
+o1-mini 87.87 98.11 80.52 89.64 81.29 88.25 85.11
+o3-mini 87.43 96.88 89.89 91.55 91.26 88.26 89.07
+Gemini-2.0-Flash 80.41 90.48 81.93 88.13 79.98 85.23 83.14
+Llama-4 87.73 94.71 81.79 89.85 85.69 89.07 86.43
+3.1 Overview of Benchmark Performance
+We begin our evaluation with a comprehensive comparison of overall accuracy across the benchmark tasks, including MATH, GSM8K, and four mathematical subsets from the MMLU corpus(MAA, MFL, MCM, MHM).
+Table 3 summarizes the accuracy score (ACC), clarity score (CS)and reasoning scores (RS), highlighting each model’s performance and robustness across these domains.
+Overall, the o3-mini model achieved the highest average accuracy, outperforming all baselines with an average of 97.06%.
+It showed particularly strong performance in the Math (96.56%), MAA (96.00%), and MCM (99.00%) subsets.
+Similarly, DeepSeek-R1 and o1 demonstrated competitive accuracy, achieving 95.75% and 96.91% respectively.
+When evaluating reasoning score, o3-mini again led the comparison with an average score of 98.45%, closely followed by DeepSeek-R1 (96.08%) and o1 (95.97%).
+These models consistently showed high reasoning coherence and correctness across all mathematical domains.
+Importantly, we include representative examples of both correct and incorrect responses in Appendix C.
+In terms of clarity score, which evaluates the explanation’s readability and logical structure, the results followed a similar trend.
+o1 and o1-mini scored highest with averages of 88.44% and 85.11%, respectively.
+Notably, GPT-4o, although not optimized explicitly for math reasoning, maintained competitive scores in clarity (87.33%) and reasoning (94.32%).
+Notably, the distilled variant DeepSeek-1.5B exhibited markedly lower scores across all three metrics, especially in reasoning (77.64%) and clarity (73.62%), emphasizing the performance compromises introduced by model compression.
+Appendix D provides a detailed comparative breakdown of each model’s performance.
+In summary, while all evaluated models show proficiency, our findings reveal significant variance in cross-domain generalization.
+The superior performance of models like DeepSeek-R1 and o1 across diverse mathematical tasks positions them as current leaders in LLM-based mathematical reasoning, while others specially deepseek’s distilled variant, despite it’s scale, remain constrained by limitations in symbolic abstraction and formal manipulation.
+3.2 Cognition Profile Across Reasoning Dimensions
+To better understand the model-wise variation in mathematical reasoning behavior, we visualize the cognition profiles of each evaluated LLM using radar plots (Figure 3).
+Each subplot corresponds to a single model, showing its relative performance across six domains: MATH, GSM8K, MAA, MFL, MCM, and MHM.
+The colored lines represent the evaluation dimensions—Accuracy, Reasoning, and Clarity—as measured by the DeepEval framework [Ip and Vongthongsri, 2025].
+The Figure 3 demonstrates, o3-mini demonstrates consistent and high performance across all dimensions, with particularly dominant accuracy and reasoning scores, forming a near-regular and saturated polygon.
+This consistency highlights the effectiveness of its architecture in capturing logical structure and factual correctness.
+Figure 3: Radar plots of LLM performance across six mathematical domains.
+Each sub-plot corresponds to a model.
+Lines represent three evaluation metrics: Accuracy, Reasoning, and Clarity.
+This visualization highlights the strengths and weaknesses of each model across different reasoning dimensions and content types.
+In contrast, the DeepSeek-1.5B (distilled) model displays notable asymmetry in its radar shape, indicating uneven performance across both domains and evaluation metrics.
+Specifically, the reduced reasoning and clarity scores in the MAA and MFL domains suggest limitations in handling abstract and symbolic reasoning, a likely result of aggressive distillation.
+On the other hand, GPT-4o model shows strong reasoning capability, especially in GSM8K and MFL, though with relatively lower accuracy in the MATH domain.
+Its clarity curve remains stable, affirming its general-purpose alignment rather than domain-specific tuning.
+DeepSeek-R1 and o1 both display high and balanced performance, especially in the reasoning dimension, validating their suitability for high-precision mathematical applications.
+Meanwhile, Llama-4 and Gemini 2.0 Flash show strong clarity but slightly lower reasoning compared to the top-tier models.
+Overall, these cognition profiles emphasize that high performance in mathematical reasoning tasks depends not just on accuracy but also on the model’s ability to maintain logical flow and human-readable clarity across domains.
+The radar plots provide an interpretable and multi-dimensional view into how different models prioritize or trade off among these cognitive capabilities.
+3.3 Accuracy–Latency Trade-off Analysis
+Figure 4: Mean accuracy (%) versus average response time (seconds) for each model.
+The bar height indicates average accuracy across reasoning tasks, while the color gradient and labels represent the average response latency.
+Efficient response times are crucial in the deployment of LLMs, as they directly influence user experience and the practicality of real-time applications.
+Factors affecting inference speed include model size, computational complexity, hardware capabilities, and optimization techniques.
+To complement our evaluation of cognitive performance, we investigate the trade-off between reasoning accuracy and computational efficiency across models (Figure 4).
+Each bar represents a model’s average accuracy across our mathematical reasoning benchmarks, while the color gradient and numerical labels denote the average response time (in seconds), measured using consistent inference settings.
+Among the models, o3-mini achieves an optimal balance, with a high accuracy of 97.1% and a relatively low response time of 6.8 seconds.
+This performance places it at the efficiency frontier, demonstrating both computational effectiveness and strong reasoning capabilities.
+Similarly, o1-mini maintains respectable accuracy (88.9%) with an even lower latency of 5.3 seconds, making it highly suitable for real-time or resource-constrained deployments.
+Furthermore, Gemini 2.0 Flash stands out for its minimal latency of just 4.2 seconds while achieving a solid 91.8% accuracy.
+This profile makes it particularly attractive for latency-sensitive applications.
+In contrast, models like o1 and DeepSeek-R1 deliver exceptional accuracy scores (96.9% and 95.8%, respectively) but exhibit significantly longer response times (15.0 and 81.05 seconds), which may present challenges in time-critical settings.
+The DeepSeek-1.5B model highlights the trade-off at the other end of the spectrum, with fast inference (5.0 seconds) but substantially lower accuracy, reflecting the compromises often seen in smaller or distilled architectures.
+Overall, this comparison underscores a clear trade-off between accuracy and inference speed.
+While some models prioritize correctness at the cost of latency, others offer rapid responses with modest performance degradation.
+4 Discussion
+In this study, we conducted a comprehensive evaluation of various LLMs to assess their mathematical reasoning capabilities across multiple benchmarks [Cobbe et al., 2021, Hendrycks et al., 2020, 2021].Our empirical analysis reveals that o3-mini outperformed all evaluated models in terms of overall accuracy and reasoning quality across a range of mathematical domains.
+It demonstrated particular strength in complex and abstract areas, such as Abstract Algebra and Multistep Calculation, while maintaining relatively low latency.
+In comparison, DeepSeek-R1 and o1 also exhibited strong reasoning capabilities, especially in structured arithmetic domains, but their performance was slightly lower than o3-mini in both accuracy and efficiency.
+These results confirm the existence of a trade-off between computational efficiency and reasoning performance, which must be carefully considered when designing LLMs for downstream applications.
+Furthermore, our results validate the advanced design of DeepSeek-R1, which employs GRPO—a reinforcement learning-based optimization approach—instead of conventional supervised finetuning [Evstafev, 2025, Liu et al., 2024a, Neha and Bhati, 2025].
+GRPO allows the model to evaluate its outputs relative to a group of sampled peers, eliminating the need for external reward models [Liu et al., 2024a, b].
+This strategy achieves strong performance, particularly on GSM8K and logic-based tasks, it remains slightly behind o3-mini in accuracy and reasoning consistency across all domains.
+These comparisons suggest that reinforcement learning with GRPO enhances model generalization, but domain-specific prompting strategies—like those used in o3-mini—can still offer an edge in specialized reasoning tasks.
+Moreover, proof-based reasoning, a crucial component of higher-level mathematics, demands meta-cognitive processes that extend beyond pattern recognition, including recursive logic application and theorem proving.
+The limitations observed in DeepSeek-R1 to generalize across abstract mathematical problems indicate that a more versatile reasoning approach may be needed to complement GRPO-based training.
+In contrast, CoT reasoning demonstrates strong performance in specific mathematical fields by leveraging step-by-step logical breakdowns to enhance problem-solving accuracy.
+Thus, while CoT provides a deeper, structured reasoning framework [Boye and Moell, 2025, Li et al., 2024a], it may struggle to satisfy the generalization requirements necessary for handling diverse mathematical challenges.
+The trade-off between GRPO’s structured logic refinement and CoT’s sequential deductive approach suggests that a hybrid reasoning mechanism may be essential for developing a more versatile and mathematically proficient LLM.
+In summary, our findings align with prior research [Ahn et al., 2024a, Akella, 2024, Duan, 2025, Evstafev, 2025, Jiang et al., 2025, Li et al., 2024b, 2025, Mercer et al., 2025, Mirzadeh et al., 2024, Peng et al., 2024, Shao et al., 2024, Xin et al., 2024, Zhong et al., 2024, Zhou et al., 2024, Zhuang et al., 2024], further confirming that the recent invented model DeepSeek-R1 demonstrates strong mathematical reasoning compared to other LLMs.
+However, our research extends beyond previous work by providing a granular, field-specific analysis of LLM performance across different domains.
+We delve into the nuances of logical, arithmetic, and abstract reasoning capabilities, offering a holistic understanding of how different LLMs excel or struggle across mathematical disciplines.
+Despite its valuable contributions, our study has several limitations: 1).
+Limited Model Scope: We did not test other cutting-edge models, such as Qwen.
+2).
+Restricted Benchmark Diversity.
+More importantly, future research should expand the scope of evaluated LLMs and incorporate additional benchmarks to gain a broader understanding of model strengths and weaknesses across varied mathematical disciplines.
+Notably, integrating GRPO with peer CoTs could allow models to leverage both structured logical optimization and stepwise deductive reasoning.
+This integration could lead to LLM swarms, where multiple reasoning frameworks synergize to enhance efficiency and accuracy.
+Meanwhile, our study reaffirms that distillation can significantly impair reasoning capabilities.
+Future efforts should explore adaptive distillation methods that preserve deep reasoning pathways while optimizing for computational efficiency.
+5 Conclusion
+Overall, our study presents a comprehensive comparative analysis of LLM performance in mathematical reasoning, identifying key strengths and weaknesses across multiple domains.
+We highlight the effectiveness of GRPO in structured problem-solving, demonstrating its ability to enhance logical consistency and rule-based reasoning.
+However, we also uncover its limitations in handling abstract mathematical concepts, where deep symbolic manipulation and theorem-based reasoning remain challenging for GRPO-trained models.
+Furthermore, our findings underscore the trade-offs associated with model distillation, revealing its potential drawbacks in preserving mathematical reasoning capabilities.
+Beyond performance evaluation, this study paves the way for future advancements in LLM-driven mathematical intelligence by outlining potential enhancements in reasoning frameworks, including hybrid models that integrate RL with structured step-by-step inference methods.
+By addressing these challenges, our work contributes to the development of next-generation AI systems capable of handling complex mathematical reasoning with greater accuracy, efficiency, and adaptability.
+6 Acknowledgement
+The authors gratefully acknowledge the support of High Performance Computing Services for providing the computational resources that contributed to the results presented in this publication/report.
+Appendix A Appendix A
+The table 2 summarizes the benchmark datasets used for evaluating the mathematical reasoning abilities of LLMs.
+The MATH dataset consists of free-form numeric problems drawn from math competitions, covering advanced topics like algebra, geometry, number theory, and combinatorics.
+GSM8K focuses on grade school arithmetic through structured word problems.
+The MMLU subsets include four college-level and high school-level categories—College Mathematics, Abstract Algebra, Formal Logic, and High School Mathematics—each formatted as multiple-choice questions and covering domains such as discrete math, algebraic structures, and logical reasoning.
+MATH is derived from high-level mathematics competitions, including the AMC 10, AMC 12, and AIME [Hendrycks et al., 2021].
+These problems are inherently more complex than standard K-12 mathematics tasks, requiring the application of non-trivial problem-solving strategies, logical inference, and domain-specific heuristics rather than direct formulaic computations.
+Meanwhile, MATH is widely regarded as a difficult benchmark, with LLMs achieving accuracy rates between 3.0% and 6.9% [Hendrycks et al., 2021].
+Despite these low overall scores, LLMs demonstrate some degree of mathematical competency, as evidenced by: Up to 15% accuracy on the easiest difficulty level [Hendrycks et al., 2021].
+The ability to generate step-by-step solutions that, while sometimes incorrect, remain coherent and contextually relevant.
+To gauge the dataset’s difficulty, human performance was also considered.
+For instance, a computer science Ph.D.
+student with no particular affinity for mathematics achieved approximately 40% accuracy.
+In addition, a three-time IMO (International Mathematical Olympiad) gold medalist attained 90% accuracy.
+Thus, these findings indicate that MATH is challenging for both human solvers and LLMs, making it an invaluable dataset for assessing and advancing mathematical reasoning capabilities in both LLMs and human problem solvers.
+GSM8K, introduced in 2021, is a benchmark dataset consisting of 8,500 high-quality grade school-level math problems  [Cobbe et al., 2021].
+Designed to incorporate high linguistic diversity while relying on fundamental mathematical concepts, it presents a unique challenge for state-of-the-art LLMs.
+Although the underlying math is relatively simple, the diverse problem formulations create significant hurdles, preventing many models from achieving consistently high accuracy.
+GSM8K is structured into 7,500 training problems and 1,320 testing problems, all carefully crafted by expert human problem writers.
+The problems primarily involve elementary arithmetic operations and typically require 2 to 8 logical steps to reach a solution.
+This dataset is widely used to evaluate logical reasoning and mathematical proficiency in LLMs and serves as a benchmark for various assessments, including the LLM Leaderboard.
+Notably, while some GSM8K problems are conceptually straightforward, they can still be challenging for even the most advanced LLMs, often exhibiting high variability in responses when the same problem is presented in slightly different ways.
+This highlights the ongoing difficulty of achieving robust mathematical reasoning in language models.
+In this study, we selected the 1.32K mathematical test set from GSM8K to systematically evaluate and compare the mathematical reasoning capabilities of each LLM.
+The MMLU dataset is a comprehensive benchmark comprising multiple-choice questions from a diverse range of academic disciplines.
+Spanning 57 distinct tasks, it covers subjects in the humanities, social sciences, hard sciences, and mathematical reasoning, reflecting the breadth of knowledge that is essential for various fields of study [Hendrycks et al., 2020].
+MMLU’s questions were manually curated by graduate and undergraduate students from publicly available sources, including practice questions from standardized exams such as the Graduate Record Examination (GRE) and the United States Medical Licensing Examination (USMLE).
+Additionally, it features questions designed for undergraduate courses and readers of Oxford University Press books.
+Notably, its mathematical reasoning component encompasses multiple subfields, categorized into areas such as "Abstract Algebra," "College Mathematics," "Formal Logic," and "High School Mathematics." For instance, questions in the "Abstract Algebra" category originate from professional mathematical practice, while the "High School Mathematics" category includes problems akin to those found in standard high school exams.
+This broad and structured dataset serves as a critical benchmark for evaluating the reasoning and problem-solving capabilities of LLMs.
 Table 2: Overview of Benchmark Data subsets Used for Evaluation
@@
 High School Mathematics Multiple-choice General Mathematics 270
-3.2.1 MATH
-MATH is a comprehensive benchmark dataset designed to evaluate mathematical reasoning across a wide range of problem-solving scenarios [40].
-Unlike conventional mathematics problem datasets, which primarily assess algorithmic proficiency, MATH focuses on advanced reasoning and heuristic-driven problem-solving.
-Specifically, MATH is derived from high-level mathematics competitions, including the AMC 10, AMC 12, and AIME [40].
-These problems are inherently more complex than standard K-12 mathematics tasks, requiring the application of non-trivial problem-solving strategies, logical inference, and domain-specific heuristics rather than direct formulaic computations.
-Meanwhile, MATH is widely regarded as a difficult benchmark, with LLMs achieving accuracy rates between 3.0% and 6.9% [40].
-Despite these low overall scores, LLMs demonstrate some degree of mathematical competency, as evidenced by: Up to 15% accuracy on the easiest difficulty level [40].
-The ability to generate step-by-step solutions that, while sometimes incorrect, remain coherent and contextually relevant.
-To gauge the dataset’s difficulty, human performance was also considered.
-For instance, a computer science Ph.D.
-student with no particular affinity for mathematics achieved approximately 40% accuracy.
-In addition, a three-time IMO (International Mathematical Olympiad) gold medalist attained 90% accuracy.
-Thus, these findings indicate that MATH is challenging for both human solvers and LLMs, making it an invaluable dataset for assessing and advancing mathematical reasoning capabilities in both LLMs and human problem solvers.
-3.2.2 GSM8K
-GSM8K, introduced in 2021, is a benchmark dataset consisting of 8,500 high-quality grade school-level math problems  [16].
-Designed to incorporate high linguistic diversity while relying on fundamental mathematical concepts, it presents a unique challenge for state-of-the-art LLMs.
-Although the underlying math is relatively simple, the diverse problem formulations create significant hurdles, preventing many models from achieving consistently high accuracy.
-GSM8K is structured into 7,500 training problems and 1,320 testing problems, all carefully crafted by expert human problem writers.
-The problems primarily involve elementary arithmetic operations and typically require 2 to 8 logical steps to reach a solution.
-This dataset is widely used to evaluate logical reasoning and mathematical proficiency in LLMs and serves as a benchmark for various assessments, including the LLM Leaderboard.
-Notably, while some GSM8K problems are conceptually straightforward, they can still be challenging for even the most advanced LLMs, often exhibiting high variability in responses when the same problem is presented in slightly different ways.
-This highlights the ongoing difficulty of achieving robust mathematical reasoning in language models.
-In this study, we selected the 1.32K mathematical test set from GSM8K to systematically evaluate and compare the mathematical reasoning capabilities of each LLM.
-3.2.3 MMLU
-The MMLU dataset is a comprehensive benchmark comprising multiple-choice questions from a diverse range of academic disciplines.
-Spanning 57 distinct tasks, it covers subjects in the humanities, social sciences, hard sciences, and mathematical reasoning, reflecting the breadth of knowledge that is essential for various fields of study [39].
-MMLU’s questions were manually curated by graduate and undergraduate students from publicly available sources, including practice questions from standardized exams such as the Graduate Record Examination (GRE) and the United States Medical Licensing Examination (USMLE).
-Additionally, it features questions designed for undergraduate courses and readers of Oxford University Press books.
-Notably, its mathematical reasoning component encompasses multiple subfields, categorized into areas such as "Abstract Algebra," "College Mathematics," "Formal Logic," and "High School Mathematics." For instance, questions in the "Abstract Algebra" category originate from professional mathematical practice, while the "High School Mathematics" category includes problems akin to those found in standard high school exams.
-In total, MMLU comprises 15,908 questions, divided into three subsets: a).
-A few-shot development set containing five questions per subject.
-b).
-A validation set with 1,540 questions, used for selecting hyperparameters.
-c).
-A test set with 14,079 questions, ensuring a rigorous evaluation Each subject within MMLU includes a minimum of 100 test questions, making it a more extensive and challenging assessment than most standard exams.
-This broad and structured dataset serves as a critical benchmark for evaluating the reasoning and problem-solving capabilities of LLMs.
-3.3 Evaluate LLMs in Mathematical Reasoning
-As discussed previously, multiple LLMs have demonstrated remarkable abilities in mathematical reasoning [105].
-To systematically evaluate the mathematical reasoning capabilities of various LLMs, we conducted experiments using API-based access to a diverse set of models.
-The selected models include Google’s Gemini-2.0-flash, OpenAI’s GPT-4o, o1-mini, o1, and o3-mini, and DeepSeek’s DeepSeek-1.5B and DeepSeek-R1.
-These models were accessed through their respective API services using secure API keys from Google, OpenAI, DeepSeek, and Openrouter.
-3.3.1 Prompting Strategy
-The effectiveness of LLMs in mathematical reasoning is highly dependent on the design of input prompts [5, 10, 25].
-To ensure a fair and reproducible evaluation, we employed a structured prompting strategy focusing on zero-shot prompting initially.
-Our methodology was designed to assess models’ inherent problem-solving abilities without additional external guidance [87].
-Overall, for all evaluations, we incorporated a zero-shot prompting approach, where models were presented with mathematical problems without any prior exemplars or demonstrations.
-Zero-shot prompting is particularly valuable in assessing a model’s ability to generalize mathematical knowledge from its pretraining corpus and apply learned concepts to novel problems [87, 98].
-By not providing explicit examples, this method evaluates how well a model can infer the appropriate solution methodology based solely on its prior learning.
+Appendix B Appendix B
+[…]
 Figure 5: Templates used to evaluate mathematical reasoning in LLMs.
 The multiple-choice format (a) ensures structured decision-making with predefined options, while the open-ended format (b) assesses free-form problem-solving capabilities.
-[…]
-For multiple-choice questions, we provided the set of answer choices and instructed the models to select the correct response from the given options, ensuring a structured decision-making process [68].
-The prompt included clear formatting guidelines to standardize the response format across different models.
-In contrast, for open-ended problems that required free-form numeric or symbolic responses, we presented the question directly without predefined answer choices, allowing the models to generate responses based on their internal reasoning capabilities [81].
-This approach ensured that multiple-choice tasks evaluated the models’ ability to differentiate between structured options, while open-ended problems tested their capability to derive solutions independently.
-3.3.2 Sampling Settings
-To ensure consistency and reliability in model responses, we employed a structured sampling strategy using temperature, top-k sampling, and top-p (nucleus) sampling [11, 70], which control randomness and diversity in token selection.
-To ensure a standardized evaluation across different LLMs, we utilized API access through OpenRouter for DeepSeek models and OpenAI API for GPT-4o, o1-mini, o1, and o3-mini, while maintaining their respective default generation settings.
-All models, including GPT-4o and o-series models, Gemini and DeepSeek use a default temperature of $T=1.0$ and an unrestricted probability sampling strategy (top-p = 1.0) [70], allowing the model to consider the full probability distribution for token selection.
-Moreover, to ensure smooth execution of API requests while avoiding rate limits imposed by different providers, we implemented an adaptive time delay between successive API calls.
-This approach prevented throttling issues while maintaining efficiency in large-scale evaluation.
-All queries were formatted in a standardized manner to ensure fair comparisons across models.
-3.3.3 Evaluation Metric
-To assess the correctness of model-generated responses in mathematical tasks, we employ the Exact Match metric as our primary evaluation method [78].
-Given that each question in our benchmark dataset has a single correct answer and the model produces a response per query, Exact Match ensures a rigorous evaluation by comparing the extracted answer to the ground truth.
-[…]
-Let $\hat{y}_{i}$ represent the extracted answer from the model’s output for the $i^{th}$ question, and let $y_{i}$ be the corresponding ground truth answer. The Exact Match accuracy is computed as:
-[…]
-where:
-  • $N$ is the total number of evaluated questions.
-  • $\mathbb{1}(\cdot)$ is the indicator function, returning $1$ if the extracted model response matches the ground truth after preprocessing, and 0 otherwise.
-  • $\text{normalize}(\cdot)$ is a function that standardizes formatting, trims spaces, and normalizes numerical values.
-Furthermore, error-handling mechanisms were incorporated to detect hallucinated responses [71, 79, 100], incomplete computations, or missing boxed outputs, with automated validation and manual review of flagged cases ensuring robust evaluation.
-Additionally, since some model responses exceeded token limits and resulted in truncated outputs, detection mechanisms were implemented to flag incomplete solutions, allowing either exclusion from evaluation or manual review when necessary.
-Table 2: Comparison of ground truth and model-generated parsed answers
-Reference answer Parsed Model answer Correctness
-p - q [’p - q’] True
-90$\circ$ [’90’] True
-4 [’5’] False
-[…]
-6 - 5i [’6 - 5i’] True
-Table 2 presents the evaluation of model-generated responses against ground truth values.
-The Reference Answer represents the ground truth, while the Parsed Model Answer contains the extracted response from the model’s output.
-The Correctness column indicates whether the model’s response exactly matches the ground truth after parsing, ensuring a strict evaluation of mathematical accuracy.
-Discrepancies, such as format mismatches or numerical errors, directly impact correctness scores, highlighting the model’s precision in solving mathematical problems.
-By integrating automated parsing, formatting corrections, and evaluation metrics, our post-processing pipeline ensured reliable and unbiased assessment of mathematical reasoning capabilities across different LLM architectures.
-4 Results
-In this section, we conducted a quantitative and qualitative evaluation of the state-of-the-art LLMs mentioned earlier, using multiple public benchmark datasets for mathematical reasoning.
-As an initial assessment, we selected a subset of three widely recognized datasets [16, 39, 62], as discussed in Section 3, to systematically evaluate the performance of various cutting-edge models.
-Overall, our evaluation includes DeepSeek-R1 [56], DeepSeek’s most advanced reasoning model, along with its distilled variant DeepSeek-1.5B [37], which is derived from the R1 model.
-Additionally, we assessed Google’s latest Gemini 2.0 Flash [45] and four OpenAI models: GPT-4o, o1-mini, o1, and o3-mini [6, 44, 103].
-Notably, while all these models are designed with reasoning capabilities, GPT-4o is the only one not explicitly optimized for providing reasoning-based CoT responses.
-These selected LLMs represent the most advanced AI systems available today, continuously evolving and competing across various mathematical reasoning benchmarks.
-4.1 Evaluation of LLMs using MATH
-As illustrated in Table 3, the performance of various LLMs on the MATH benchmark dataset [40] exhibits significant variability, underscoring differences in their mathematical reasoning capabilities.
-In particular, GPT-4o [44] exhibited a significantly lower accuracy of 64.88%, lagging behind its peer models.
-This limitation may be attributed to deficiencies in complex reasoning capabilities or the inherent difficulty of competition-level mathematical problems in the dataset.
-In contrast, OpenAI’s o1 model achieved the highest accuracy (93.12%), closely followed by DeepSeek-R1 (90.45%), with both models surpassing the 90% correctness threshold—a benchmark that no other model reached.
-Other models demonstrated varying levels of performance: Gemini 2.0 Flash, 85.87%; o1-mini, 88.93%; o3-mini, 82.06%.
-However, a key observation is that the distilled variant of DeepSeek-R1 performed significantly worse, further reinforcing the hypothesis that model distillation can impair mathematical reasoning capabilities by compromising critical reasoning pathways in favor of improved computational efficiency.
-Additionally, the quantitative comparisons across multiple LLMs, as presented in Figure 3, reveal consistent performance patterns across OpenAI’s models, while DeepSeek-R1 consistently outperformed its distilled variant.
-These findings highlight the trade-offs between model size and reasoning ability, as demonstrated by Fu et al.
-[33], underscoring the need for optimization techniques that enhance computational efficiency without sacrificing mathematical problem-solving proficiency.
-Table 3: Performance comparison of various models on MATH in terms of accuracy.
-Publisher Model Name Accuracy (%)
-Google Gemini 2.0 Flash 85.87
-OpenAI GPT-4o 64.88
-o1-mini 88.93
-o1 93.12
-o3-mini 82.06
-DeepSeek DeepSeek-1.5B 65.64
-DeepSeek-R1 90.45
-Figure 3: Performance comparison of various LLMs in the MATH benchmark.
-The accuracy of each model is displayed above the corresponding bar, highlighting differences in mathematical problem-solving capabilities across model publishers.
-Furthermore, Figure 4 provides an in-depth comparison of the prompt responses and mathematical reasoning of DeepSeek-R1, o1, and Gemini 2.0 Flash.
-Interestingly, Gemini 2.0 Flash failed to provide a correct answer, whereas DeepSeek-R1 and o1 demonstrated more structured and logically coherent reasoning processes.
-This further illustrates the varying degrees of mathematical reasoning proficiency across models [103] and the importance of structured reasoning frameworks in achieving high accuracy.
-Moreover, Figure 4 illustrates that DeepSeek-R1 and o1 followed nearly identical reasoning steps, correctly assigning coordinates, computing midpoints and centroids, and using determinant-based area calculations to arrive at the correct answer, 8.
-However, Gemini 2.0 Flash deviated from the correct reasoning path by making errors in its calculations.
-Specifically, despite correctly identifying some intermediate relationships, its handling of area scaling and altitude ratio calculations led to an incorrect final answer of 4 instead of, correct response of 8.
-This demonstrates a key limitation in Gemini 2.0 Flash’s numerical precision and reasoning accuracy for complex geometric problems compared with its peers.
-Notably, the strong performance of DeepSeek-R1 in Figure 4, which relies heavily on RL methodologies [29], highlights the effectiveness of iterative, feedback-driven training paradigms in tackling mathematically intricate tasks.
-A crucial component in its training is Group Relative Policy Optimization (GRPO) [76], a more efficient RL approach than traditional methods.
-GRPO enables DeepSeek-R1 to excel in complex domains such as mathematics, science, and coding, reinforcing the potential of advanced learning techniques in enhancing model performance [37].
-These findings underscore the importance of advanced reasoning mechanisms in solving competition-level mathematical problems, suggesting that RL frameworks-exemplified by DeepSeek-R1-represent a promising direction for enhancing the mathematical reasoning capabilities of next-generation LLMs.
+Appendix C Appendix C
+To provide a deeper understanding of the qualitative behavior of the evaluated models, we include representative examples of both correct and incorrect responses in this section.
+These examples span problems from GSM8K, MMLU, and MATH benchmarks and highlight a range of model behaviors including accurate multi-step reasoning, partial computations, and common logical problems.
+Figure 6: Comparative evaluation of different LLMs on an Arithmetic problem from the GSM8K benchmark.
+The figure showcases step-by-step reasoning from three different models and highlights correctness.
+The question is shown in red, while the extracted answers are boxed.
+Green-highlighted sections indicate correct responses, whereas red-highlighted sections denote incorrect outputs.
+This comparison illustrates differences in mathematical reasoning and accuracy across models
+Figure 7: Comparative evaluation of different LLMs on a Formal Logic problem from the MMLU benchmark.
+The figure showcases step-by-step reasoning from three different models and highlights correctness.
+The question is shown in red, while the extracted answers are boxed.
+Green-highlighted sections indicate correct responses, whereas red-highlighted sections denote incorrect outputs.
+This comparison illustrates differences in mathematical reasoning and accuracy across models.
 Figure 8: Comparative evaluation of different LLMs on a Geometric problem from the MATH benchmark.
@@
 This comparison illustrates differences in mathematical reasoning and accuracy across models.
-4.2 Evaluation of LLMs using GSM8K
-As discussed in Section 3, the GSM8K benchmark dataset [16] comprises approximately 1.32k test samples designed to assess math reasoning in LLMs [16].
-The zero-shot performance of various models is summarized in Table 4, presenting valuable insights into their mathematical reasoning capabilities.
-In particular, the results revealed that DeepSeek-R1 [37] performed on par with OpenAI’s o1, with both models achieving the highest accuracy of 96.13%, surpassing all other peers in this evaluation.
-Additionally, several other models demonstrated strong performance: Gemini 2.0 Flash, 95.53%; GPT-4o, 95.58%; o1-mini, 95.91%; o3-mini, 95.83% However, a notable performance gap was observed in DeepSeek-R1’s distilled variant, which achieved only 81.13% accuracy—a significant decline compared to its full-scale counterpart.
-This performance disparity underscores the trade-offs associated with model distillation, where reducing the number of parameters for computational efficiency may inadvertently weaken logical reasoning and problem-solving capabilities.
-While distillation enables smaller, faster models, our findings suggest that excessive parameter reduction can compromise mathematical reasoning proficiency, potentially limiting their effectiveness in complex problem-solving tasks.
-Similarly, Figure 5 demonstrates that DeepSeek-R1 and o1 outperformed other peer LLMs.
-Notably, there are no significant differences across all LLMs released via OpenAI.
-These results emphasize the delicate balance between model size and reasoning ability, highlighting the need for advancing distillation techniques that preserve core reasoning structures while maintaining computational efficiency.
-Table 4: Performance comparison of various models on GSM8K in terms of accuracy.
-Publisher Model Name Accuracy (%)
-Google Gemini 2.0 Flash 95.53
-OpenAI GPT-4o 95.98
-o1-mini 95.91
-o1 96.13
-o3-mini 95.83
-DeepSeek DeepSeek-1.5B 81.12
-DeepSeek-R1 96.13
-Moreover, the qualitative comparisons across multiple LLMs, as illustrated in Figure 6, reveal a noteworthy disparity in reasoning capabilities among OpenAI’s o1, DeepSeek-R1, and Google’s Gemini 2.0 Flash.
-The analysis demonstrates that DeepSeek-R1 exhibits superior performance in mathematical reasoning compared to its counterparts.
-Specifically, DeepSeek-R1 generated accurate solutions and also articulated a systematic, step-by-step process to arrive at its conclusions.
-In contrast, both OpenAI’s o1 and Gemini 2.0 Flash produced less comprehensive pathways, which resulted in erroneous answers.
-These findings, as shown in Figure 6, underscore DeepSeek-R1’s robust analytical proficiency relative to contemporary models, highlighting its advanced problem-solving capabilities in complex mathematical tasks.
-Figure 5: Performance comparison of various LLMs in the GSM8K benchmark.
+Appendix D Appendix D
+Figure 9: Performance comparison of various LLMs in the GSM8K and MATH benchmark.
 The accuracy of each model is displayed above the corresponding bar, highlighting differences in mathematical problem-solving capabilities across model publishers.
-Figure 6: Comparative evaluation of different LLMs on an Arithmetic problem from the GSM8K benchmark.
-The figure showcases step-by-step reasoning from three different models and highlights correctness.
-The question is shown in red, while the extracted answers are boxed.
-Green-highlighted sections indicate correct responses, whereas red-highlighted sections denote incorrect outputs.
-This comparison illustrates differences in mathematical reasoning and accuracy across models.
-4.3 Evaluation of LLMs using MMLU
-Leveraging the structured categorization of MMLU benchmark dataset [39], we conduct an in-depth comparative analysis of various LLMs across different mathematical domains.
-First, as shown in Table 5, DeepSeek-R1 outperformed all peer models, achieving the highest accuracy of 97.62% in Formal Logic.
-Other models, such as o1, o3-mini, and Gemini 2.0 Flash, also demonstrated strong performance, all surpassing 90% accuracy.
-However, distilled variants of DeepSeek-R1 and o1-mini failed to reach this benchmark, reinforcing the hypothesis that model distillation can degrade logical reasoning capabilities by limiting access to essential computational pathways.
-Table 5: Performance comparison of various models on MMLU in terms of accuracy.
-Problem Type Model Name Accuracy (%)
-Abstract Algebra Gemini 2.0 flash 87.00
-GPT-4o 85.00
-o1-mini 88.00
-o1 94.00
-o3-mini 96.00
-DeepSeek-1.5B 61.00
-DeepSeek-R1 94.00
-College Mathematics Gemini 2.0 flash 96.00
-GPT-4o 79.00
-o1-mini 76.00
-o1 99.00
-o3-mini 99.00
-DeepSeek-1.5B 65.00
-DeepSeek-R1 96.00
-Formal Logic Gemini 2.0 flash 90.48
-GPT-4o 82.54
-o1-mini 78.57
-o1 95.24
-o3-mini 96.03
-DeepSeek-1.5B 47.62
-DeepSeek-R1 97.62
-High School Mathematics Gemini 2.0 flash 97.77
-GPT-4o 88.15
-o1-mini 98.89
-o1 99.26
-o3-mini 98.52
-DeepSeek-1.5B 86.30
-DeepSeek-R1 97.04
-In addition, in Abstract Algebra, o3-mini outperformed its peers, achieving the highest accuracy of 96.00%.
-Abstract algebra is a particularly challenging subject, even for math graduates, as it demands a fundamental shift in mathematical thinking from traditional algebraic manipulations to analyzing complex algebraic structures such as groups, rings, and fields.
-Mastery of this subject requires a deep understanding of abstract concepts, rigorous proof construction, and advanced logical reasoning.
-The results also suggest that OpenAI’s models, particularly o3-mini, excel in abstract reasoning and symbolic manipulation, enabling them to reconstruct and process complex algebraic structures effectively.
-In contrast, the distilled variant of DeepSeek-R1 performed significantly worse, achieving only 61% accuracy, further validating the detrimental impact of parameter reduction on logical inference and abstract reasoning.
-Furthermore, in College Mathematics, both o1 and o3-mini achieved exceptional accuracy of 99.00% in College Mathematics, surpassing all other models.
-This outstanding performance can be attributed to two key factors: a).
-Superior reasoning capabilities: OpenAI’s o1 and o3-mini models are designed with enhanced mathematical reasoning frameworks, allowing them to efficiently process and solve complex college-level problems.
-b).
-Diversity of training data: The extensive dataset exposure during training may have equipped o1 and o3-mini with a broader mathematical foundation, leading to their unparalleled performance in advanced mathematical reasoning tasks.
-Meanwhile, Gemini 2.0 Flash also demonstrated strong performance, achieving a comparable accuracy of 96%, reinforcing its competence in college-level mathematics.
-However, distillation had a clear negative impact on performance, as evidenced by: 1).
-o1-mini, which achieved only 76.00% accuracy, a significant drop from its full-scale counterpart.
-2).
-DeepSeek-1.5B, which attained just 65.00% accuracy, further emphasizing the trade-off between model size and reasoning ability.
-This performance gap highlights the increased complexity of college mathematics compared to high school-level math.
-College-level coursework usually demands: 1).
-A deeper understanding of abstract mathematical concepts; 2).
-Stronger logical reasoning abilities; 3).
-Proficiency in complex proofs and real-world applications.
-Notably, the faster pace and heavier workload of college mathematics compared to high school math further contribute to its difficulty [7].
-Moreover, given the relative simplicity of High School Mathematics, o1 achieved the highest accuracy at 99.62%, slightly improving upon its results in College Mathematics.
-Other peer models also performed well: DeepSeek-R1, 97.04%; o1-mini, 98.89%; o3-mini, 98.52%; Gemini 2.0 Flash, 97.77%.
-GPT-4o showed a lower accuracy (88.15%) compared with its peers in this domain too.
-Once again, the distilled variant of DeepSeek-R1 struggled, achieving only 86.30% accuracy, further demonstrating the performance degradation caused by model compression.
-Similarly, Figure 7 presents a comprehensive comparative analysis of various LLMs across four key mathematical domains: Abstract Algebra, College Mathematics, Formal Logic, and High School Mathematics.
-Notably, OpenAI’s models demonstrate superior performance in abstract reasoning, particularly excelling in Abstract Algebra, where symbolic manipulation and deep theoretical understanding are crucial.
-This suggests that OpenAI’s models have strong capabilities in reconstructing and processing abstract mathematical structures.
-Significantly, DeepSeek-R1 outperformed its peers in Formal Logic, showcasing its exceptional logical reasoning skills and ability to handle complex rule-based problem-solving more effectively than other models.
-These findings further highlight the specialized strengths of different LLMs and enhance the need for targeted model improvements to optimize performance across various mathematical domains.
+Figure 10: Accuracy comparison of large language models (LLMs) across four mathematical MMLU sub-domains: College Mathematics, Abstract Algebra, Formal Logic, and High School Mathematics.
+This visualization highlights performance variance across both model architectures and mathematical domains, showcasing how some models (e.g., o3-mini, o1) maintain consistently high scores across all categories, while others (e.g., DeepSeek-1.5B) struggle in more abstract reasoning tasks like Formal Logic and Algebra.
 Figure 11: Comparison of LLM models across different problem types in the MMLU benchmark dataset.
-Figure 8: Comparative evaluation of different LLMs on a Formal Logic problem from the MMLU benchmark.
-The figure showcases step-by-step reasoning from three different models and highlights correctness.
-The question is shown in red, while the extracted answers are boxed.
-Green-highlighted sections indicate correct responses, whereas red-highlighted sections denote incorrect outputs.
-This comparison illustrates differences in mathematical reasoning and accuracy across models.
-Figure 8 presents a comparative evaluation of different LLMs on a Formal Logic problem from the MMLU benchmark dataset [39].
-The task involves identifying the correct conclusion from a given argument, with four answer choices labeled A, B, C, and D.
-The correct reference answer is option D, requiring models to distinguish between supporting premises and the actual conclusion.
-DeepSeek-R1 successfully identifies option D as the correct conclusion, offering a structured explanation by recognizing that the phrase "Because of this" serves as a linguistic connector rather than a part of the conclusion itself.
-In contrast, o1 and Gemini 2.0 Flash select option C, misinterpreting the logical structure of the argument and failing to differentiate between an explanatory transition and the central claim.
-The comparative analysis in Figure 8 further solidifies DeepSeek-R1’s exceptional logical reasoning skills and ability to handle complex rule-based problem-solving tasks.
-Figure 9: Comparative evaluation of different LLMs on an Abstract Algebra problem from the MMLU benchmark.
-The figure showcases step-by-step reasoning from three different models and highlights correctness.
-The question is shown in red, while the extracted answers are boxed.
-Green-highlighted sections indicate correct responses, whereas red-highlighted sections denote incorrect outputs.
-This comparison illustrates differences in mathematical reasoning and accuracy across models.
-Figure 9 illustrates a comparative evaluation of different LLMs in solving a problem related to abstract algebra, sourced from the MMLU benchmark dataset.
-The problem required reasoning about the number of elements of a given order in a group, specifically evaluating whether Statement 1 and Statement 2 hold true under group-theoretic principles.
-The task involved selecting the correct truth-value combination from four possible answer choices: A (True, True), B (False, False), C (True, False), and D (False, True).
-The reference answer, indicated in the Figure 9, is option A (True, True).
-DeepSeek-R1 correctly identified both statements as true, providing a detailed justification using the properties of cyclic subgroups, Euler’s totient function, and disjoint subgroup structures.
-Similarly, o1 also arrived at the correct answer, correctly recognizing that exceeding 8 elements of order 15 necessitates at least 16 elements due to the existence of multiple distinct subgroups.
-However, Gemini 2.0 Flash incorrectly classified Statement 2 as false, demonstrating a reasoning flaw in its approach to subgroup structures and element counting.
-This comparison in Figure 9 highlights DeepSeek-R1 and o1’s superior reasoning abilities in abstract algebra, as both models successfully applied cyclic subgroup properties and disjoint set reasoning to reach the correct conclusion.
-In summary, DeepSeek-R1 dominated in Formal Logic, while o3-mini excelled in Abstract Algebra.
-o1 and o3-mini outperformed all models in College Mathematics, likely due to their advanced reasoning frameworks and diverse training data.
-Model distillation significantly reduced performance as shown in [37], particularly in complex reasoning tasks like Abstract Algebra and College Mathematics.
-High school-level mathematics is relatively easier for LLMs, with multiple models surpassing 97% accuracy.
-These findings emphasize the trade-offs between model size, reasoning capability, and computational efficiency, highlighting the need for optimization strategies that maintain problem-solving proficiency while improving scalability.
-4.4 Statistical Analyses
-In this section, Table 6 presents a comparative evaluation of LLMs based on their mean accuracy and standard deviation across all three benchmark datasets.
-Among the evaluated models, OpenAI’s o1 achieves the highest mean accuracy (96.13%), followed closely by DeepSeek-R1 (95.21%) and o3-mini (94.57%).
-While Gemini 2.0 Flash attains a relatively high accuracy of 92.11%, it underperforms relative to the top-tier models.
-In terms of stability, as quantified by standard deviation, both o1 (2.32) and DeepSeek-R1 (2.41) exhibit the most consistent performance.
-We believe that DeepSeek-R1’s high consistency can be attributed to its GRPO framework [37], which utilizes a multi-agent collaborative architecture, effectively integrates diverse reasoning pathways.
-This synergy helps mitigate task-specific biases and enhances consistency by reducing variability.
-The structured approach allows the model to dynamically prioritize high-confidence solutions while iteratively refining less certain outputs, thereby contributing to its robustness across heterogeneous datasets.
-By contrast, o1-mini demonstrated significantly higher variability (standard deviation: 8.31), suggesting sensitivity to dataset-specific characteristics.
-Gemini 2.0 Flash (4.59) and o3-mini (5.74) exhibited moderate fluctuations in performance, likely due to limitations in their single-agent reasoning paradigms.
-GPT-4o also exhibited a significantly larger standard deviation (10.42) and was one of the two most inconsistent models alongside DeepSeek-1.5B.
-Meanwhile, DeepSeek-1.5B’s considerably lower mean accuracy (67.78%) and the highest standard deviation (12.82) reflect substantial variability in its responses, further cementing the detrimental impact of parameter reduction on mathematical reasoning.
-Table 6: Performance statistics of different LLM models across all datasets, including mean accuracy (in %) and standard deviation.
-LLM Models Gemini 2.0 Flash GPT-4o o1-mini o1 o3-mini DeepSeek-1.5B DeepSeek-R1
-Mean (%) 92.11 82.59 87.72 96.13 94.57 67.78 95.21
-Standard Deviation 4.59 10.42 8.31 2.32 5.74 12.82 2.41
-Figure 10: Performance comparison of five LLMs across multiple datasets, grouped by publisher.
-The bars represent the mean accuracy (%), while the error bars indicate the standard deviation across three benchmark datasets.
-Figure 10 presents the performance statistics of different LLMs across three datasets.
-The bar chart in Figure 10 illustrates the mean accuracy (%) of each model, with error bars representing the standard deviation of the accuracy percentages across datasets.
-The standard deviation bars indicate the variability in the performance of each model.
-Notably, o1 from OpenAI demonstrates the highest mean accuracy with a low variance, whereas DeepSeek-1.5B exhibits the largest variability, suggesting inconsistency in its accuracy across datasets.
-This visualization underscores that o1 and DeepSeek-R1 excel not only in accuracy but also in reliability for mathematical problem-solving tasks.
-4.5 Response time analysis
-Efficient response times are crucial in the deployment of LLMs, as they directly influence user experience and the practicality of real-time applications.
-Factors affecting inference speed include model size, computational complexity, hardware capabilities, and optimization techniques.
-Figure 11: Latency Comparison across different LLMs
-Notably, Figure 11 presents a comparative analysis of the average response time per query across various LLMs.
-The DeepSeek-R1 model exhibits the highest latency, averaging 81.0 seconds per response, which implies that DeepSeek-R1 is significantly slower than all other models.
-Among OpenAI’s models, o1 demonstrates a high response time of 15.0 seconds, likely due to its complex CoT process.
-GPT-4o also exhibits relatively high latency at 11.5 seconds, making it one of the three models that exceed the 10-seconds latency threshold.
-In contrast, o1-mini and o3-mini achieved faster inference speeds of 6.8 seconds and 5.3 seconds, respectively, outperforming the other two OpenAI models.
-DeepSeek-1.5B operates at 5.0 seconds, slightly outperforming OpenAI’s o3-mini models.
-Notably, Gemini 2.0 Flash exhibits the fastest response time at 4.2 seconds, making it the most efficient model in terms of inference speed.
-Importantly, understanding these response times is essential for selecting appropriate models based on application requirements, balancing the trade-offs between latency and performance.
-5 Discussion
-In this study, we conducted a comprehensive evaluation of various LLMs to assess their mathematical reasoning capabilities across multiple benchmarks [16, 39, 40].
-The evaluated LLMs include Gemini 2.0 Flash [45], GPT-4o [44], o1-mini [46], o1 [103], o3-mini [6], DeepSeek-1.5B [57], and DeepSeek-R1 [56].
-Our empirical analysis revealed that DeepSeek-R1 and o1 outperformed its peers in most mathematical domains, demonstrating strong mathematical reasoning capabilities in solving structured problems.
-However, its performance was not on par with o3-mini in some complex and abstract fields, such as Abstract Algebra.
-Additionally, our findings underscore the impact of model distillation on mathematical reasoning performance.
-Specifically, distilled variants showcased noticeable declines in accuracy, reinforcing concerns that distillation, while improving computational efficiency, may inadvertently weaken the LLM’s capability to handle intricate problem-solving tasks.
-These observations suggest that there exists a trade-off between computational efficiency and reasoning depth, which needs to be carefully managed in future model optimizations.
-First, our results further validate the advanced capabilities of DeepSeek-R1, which is trained using RL rather than Supervised Fine-Tuning (SFT) [31, 56, 69].
-DeepSeek-R1 incorporates GRPO, a specialized RL-based optimization strategy that enhances training efficiency by evaluating model actions relative to a group of sampled peers.
-Unlike traditional RL approaches that require an external reward model, GRPO eliminates the need for a separate reward function by dynamically calculating advantages based on group-based scoring [56, 57].
-This unique training paradigm delivers DeepSeek-R1 to balance cost and performance efficiently, leading to competitive accuracy in solving grade school-level mathematical problems.
-The superior performance of DeepSeek-R1 in GSM8K highlights its superiority in handling structured, multi-step arithmetic reasoning tasks.
-Notably, DeepSeek-R1 remains comparable to o1 and o3-mini, demonstrating that while GRPO enhances efficiency, it does not necessarily surpass models with specific CoT in all problem categories.
-In addition, our analysis of DeepSeek-R1’s performance across MMLU benchmarks reveals a significant variation in mathematical reasoning capabilities across different domains.
-While DeepSeek-R1 excelled in Formal Logic, it struggled in more abstract mathematical areas, such as College Mathematics and Abstract Algebra.
-In particular, this discrepancy suggests that GRPO is particularly effective for solving problems that rely on structured logical reasoning, where a group-based approach to RL can refine logical consistency and decision-making [34].
-Formal Logic problems, for instance, often involve pattern recognition and structured deduction [34], which could align well with GRPO’s framework.
-The group scoring mechanism in GRPO enables the model to internalize patterns commonly recognized by human reasoning, allowing it to perform well in problems that require rule-based logical deductions.
-However, the performance drop in Abstract Algebra and College Mathematics suggests that DeepSeek-R1 struggles with tasks requiring highly abstract conceptualization.
-Abstract Algebra, in particular, represents a fundamental shift from procedural algebra to the exploration of abstract mathematical structures such as groups, rings, and fields [99].
-Unlike rule-based logic problems, abstract algebra necessitates deep symbolic manipulation, theorem proofs, and the ability to navigate complex algebraic structures [99], which require advanced deductive reasoning beyond GRPO’s core optimization strategy.
-Furthermore, proof-based reasoning, a crucial component of higher-level mathematics, demands meta-cognitive processes that extend beyond pattern recognition, including recursive logic application and theorem proving.
-The limitations observed in DeepSeek-R1 to generalize across abstract mathematical problems indicate that a more versatile reasoning approach may be needed to complement GRPO-based training.
-In contrast, CoT reasoning demonstrates strong performance in specific mathematical fields by leveraging step-by-step logical breakdowns to enhance problem-solving accuracy.
-CoT excels in tasks that require explicit intermediate reasoning steps, which helps models navigate complex mathematical derivations more effectively.
-However, despite these advantages, CoT remains constrained by its tendency to overfit to structured problem-solving paradigms [9, 50], leading to limited adaptability when faced with problems requiring broader conceptual generalization.
-Thus, while CoT provides a deeper, structured reasoning framework, it may struggle to satisfy the generalization requirements necessary for handling diverse mathematical challenges.
-The trade-off between GRPO’s structured logic refinement and CoT’s sequential deductive approach suggests that a hybrid reasoning mechanism—integrating the strengths of both frameworks—may be essential for developing a more versatile and mathematically proficient LLM.
-In summary, our findings align with prior research [2, 4, 27, 31, 47, 52, 54, 64, 66, 72, 76, 91, 103, 105, 107], further confirming that DeepSeek-R1 demonstrates strong mathematical reasoning compared to other LLMs.
-However, our research extends beyond previous work by providing a granular, field-specific analysis of LLM performance across different domains.
-Instead of evaluating general reasoning skills, we delve into the nuances of logical, arithmetic, and abstract reasoning capabilities, offering a holistic understanding of how different LLMs excel or struggle across mathematical disciplines.
-Despite its valuable contributions, our study has several limitations: 1).
-Limited Model Scope: While our evaluation includes prominent LLMs, we did not test other cutting-edge models, (such as Llama-3.2 / Llama-3.3, Claude, Mistral) which could provide additional insights into alternative training methodologies.
-2).
-Restricted Benchmark Diversity: Our assessment relies on three primary benchmark datasets, which, while comprehensive, may not fully capture the breadth of mathematical reasoning challenges in real-world applications.
-More importantly, future research should expand the scope of evaluated LLMs and incorporate additional benchmarks to gain a broader understanding of model strengths and weaknesses across varied mathematical disciplines.
-Specifically, integrating GRPO with peer CoTs could allow models to leverage both structured logical optimization and stepwise deductive reasoning.
-This integration could lead to LLM swarms, where multiple reasoning frameworks synergize to enhance efficiency and accuracy.
-Meanwhile, our study reaffirms that distillation can significantly impair reasoning capabilities.
-Future efforts should explore adaptive distillation methods that preserve deep reasoning pathways while optimizing for computational efficiency.
-Beyond current CoT frameworks [14, 53, 88, 89, 95], investigators could provide new insights into advanced reasoning frameworks, such as Chain of Draft (CoD) [93].
-Those into how LLMs handle interconnected logical dependencies will improve abstract mathematical problem-solving [15].
-Future models can strike a balance between efficiency, depth, and mathematical intelligence by advancing multi-framework LLMs and refining distillation strategies.
-Meanwhile, Xia et al.
-[88] introduced ReasonEval, a novel methodology for evaluating the quality of reasoning steps in mathematical tasks, emphasizing the need to assess both validity and redundancy in intermediate steps.
-Additionally, Frieder et al.
-[32] explored the mathematical capabilities of ChatGPT and GPT-4, demonstrating their effectiveness in querying mathematical facts and analyzing their performance across varying levels of mathematical complexity.
-These studies lay the foundation for more comprehensive and rigorous evaluations in the future, advancing our understanding of LLMs in mathematical reasoning.
-6 Conclusion
-Overall, our study presents a comprehensive comparative analysis of LLM performance in mathematical reasoning, identifying key strengths and weaknesses across multiple domains.
-We highlight the effectiveness of GRPO in structured problem-solving, demonstrating its ability to enhance logical consistency and rule-based reasoning.
-However, we also uncover its limitations in handling abstract mathematical concepts, where deep symbolic manipulation and theorem-based reasoning remain challenging for GRPO-trained models.
-Furthermore, our findings underscore the trade-offs associated with model distillation, revealing its potential drawbacks in preserving mathematical reasoning capabilities.
-As a result, we propose new strategies for refining distillation techniques, ensuring that compressed models maintain robust problem-solving proficiency without sacrificing computational efficiency.
-Beyond performance evaluation, this study paves the way for future advancements in LLM-driven mathematical intelligence by outlining potential enhancements in reasoning frameworks, including hybrid models that integrate RL with structured step-by-step inference methods.
-By addressing these challenges, our work contributes to the development of next-generation AI systems capable of handling complex mathematical reasoning with greater accuracy, efficiency, and adaptability.
-7 Acknowledgement
-The authors acknowledge the support of the Augusta University High Performance Computing Services (AUHPCS) for providing computational resources contributing to the results presented in this publication/report.